### v4

# Three-term Machin-type formulae

Abstract.
We shall show that there exist only finitely many nondegenerate three-term Machin-type formulae and give explicit upper bounds for the sizes of variables.
Key words and phrases:
Exponential diophantine equation, Ramanujan-Nagell equation, Machin-type formula, Linear forms of three logarithms
1991 Mathematics Subject Classification:
11D61, 11Y60, 33B10
1. Introduction
Machin’s formula
(1) $4\arctan\frac{1}{5}-\arctan\frac{1}{239}=\frac{\pi}{4}$
is well known and has been used to calculate approximate values of $\pi$. Analogous formulae
(2) $\arctan\frac{1}{2}+\arctan\frac{1}{3}=\frac{\pi}{4},$
(3) $2\arctan\frac{1}{2}-\arctan\frac{1}{7}=\frac{\pi}{4}$
and
(4) $2\arctan\frac{1}{3}+\arctan\frac{1}{7}=\frac{\pi}{4},$
which are also well known, were attributed to Euler, Hutton and Hermann, respectively. But according to Tweddle [15], these formulae also seem to have been found by Machin.
Several three-term formulae such as
(5) $8\arctan\frac{1}{10}-\arctan\frac{1}{239}-4\arctan\frac{1}{515}=\frac{\pi}{4}$
due to Simson in 1723 (see [15]) and
(6) $12\arctan\frac{1}{18}+8\arctan\frac{1}{57}-5\arctan\frac{1}{239}=\frac{\pi}{4}$
due to Gauss in 1863 also have been known.
More generally, an $n$-term Machin-type formula is defined to be an identity of the form
(7) $y_{1}\arctan\frac{1}{x_{1}}+y_{2}\arctan\frac{1}{x_{2}}+\cdots+y_{n}\arctan\frac{1}{x_{n}}=\frac{r\pi}{4}$
with integers $x_{1},x_{2},\ldots,x_{n},y_{1},y_{2},\ldots,y_{n}$ and $r\neq 0$.
Theoretical studies of Machin-type formulae have begun with a series of works of Størmer, who proved that (1)-(4) are all two-term ones in 1895 [12] and gave a necessary and sufficient condition for given integers $x_{1},x_{2},\ldots,x_{n}>1$ to have a Machin-type formula (7) and $102$ three-term ones in 1896 [13]. Størmer asked for other three-term Machin-type formulae and questioned whether there exist infinitely many ones or not. Up to now the only known other nontrivial (i.e. not derived from (2)-(4)) three-term formulae are
(8) $5\arctan\frac{1}{2}+2\arctan\frac{1}{53}+\arctan\frac{1}{4443}=\frac{3\pi}{4},$
(9) $5\arctan\frac{1}{3}-2\arctan\frac{1}{53}-\arctan\frac{1}{4443}=\frac{\pi}{2},$
and
(10) $5\arctan\frac{1}{7}+4\arctan\frac{1}{53}+2\arctan\frac{1}{4443}=\frac{\pi}{4}.$
[16] attributes these formulae to Wrench [18] although these formulae cannot be found there. We note that the second and the third formulae follow from the first formula using (2) and (3) respectively.
The purpose of this paper is to answer to Størmer’s other question in negative. We shall show that there exist only finitely many three-term Machin-type formulae which do not arise from a linear combinations of three identities (2)-(4).
Størmer’s criterion is essentially as follows: For given integers $x_{1},x_{2},\ldots,x_{n}>1$, (7) holds for some integers $y_{1},y_{2},\ldots,y_{n}$ and $r\neq 0$ if and only if there exist integers $s_{i,j}(i=1,2,\ldots,n,j=1,2,\ldots,n-1)$ and Gaussian integers $\eta_{1},\eta_{2},\ldots,\eta_{n-1}$ such that
(11) $\left[\frac{x_{i}+{\sqrt{-1}}}{x_{i}-{\sqrt{-1}}}\right]=\left[\frac{\eta_{1}}{\bar{\eta}_{1}}\right]^{\pm s_{i,1}}\left[\frac{\eta_{2}}{\bar{\eta}_{2}}\right]^{\pm s_{i,2}}\cdots\left[\frac{\eta_{n-1}}{\bar{\eta}_{n-1}}\right]^{\pm s_{i,n-1}}.$
for $i=1,2,\ldots,n$.
Writing $m_{j}=\eta_{j}\bar{\eta}_{j}$ for $j=1,2,\ldots,n-1$, this condition can be reformulated as follows: there exist nonnegative integers $s_{i,j}(i=1,2,\ldots,n,j=1,2,\ldots,n)$ with $0\leq s_{i,n}\leq 1$ such that the equation
(12) $x_{i}^{2}+1=2^{s_{i,n}}m_{1}^{s_{i,1}}m_{2}^{s_{i,2}}\cdots m_{n-1}^{s_{i,n-1}}$
holds for $i=1,2,\ldots n$ and, additionally, $x_{i}\equiv\pm x_{j}\ \left(\mathrm{mod}\ m_{k}\right)$ for three indices $i,j,k$ with $x_{i}^{2}+1\equiv x_{j}^{2}+1\equiv 0\ \left(\mathrm{mod}\ m_{k}\right)$.
Thus, for given three integers $x_{1},x_{2},x_{3}>1$, there exist nonzero integers $y_{1}$, $y_{2}$, $y_{3}$, and $r$ such that $\gcd(y_{1},y_{2},y_{3})=1$ and a three-term Machin-type formula
(13) $y_{1}\arctan\frac{1}{x_{1}}+y_{2}\arctan\frac{1}{x_{2}}+y_{3}\arctan\frac{1}{x_{3}}=\frac{r\pi}{4}$
holds if and only if there exist integers $k_{i},\ell_{i}(i=1,2,3)$ and Gaussian integers $\eta_{1},\eta_{2}$ such that
(14) $\left[\frac{x_{i}+{\sqrt{-1}}}{x_{i}-{\sqrt{-1}}}\right]=\left[\frac{\eta_{1}}{\bar{\eta}_{1}}\right]^{\pm k_{i}}\left[\frac{\eta_{2}}{\bar{\eta}_{2}}\right]^{\pm\ell_{i}}$
holds for $i=1,2,3$ or, equivalently, writing $m_{j}=\eta_{j}\bar{\eta}_{j}$ for $j=1,2,\ldots,n-1$ and choosing $v_{i}\in\{0,1\}$ appropriately, the equation
(15) $x_{i}^{2}+1=2^{v_{i}}m_{1}^{k_{i}}m_{2}^{\ell_{i}}$
holds for $i=1,2,3$ and, additionally, $x_{i}\equiv\pm x_{i^{\prime}}\ \left(\mathrm{mod}\ m_{j}\right)$ for two indices $i,i^{\prime}$ with $x_{i}^{2}+1\equiv x_{i^{\prime}}^{2}+1\equiv 0\ \left(\mathrm{mod}\ m_{j}\right)$. Furthermore, there exists a nonzero integers $d$ such that $dy_{1}=\pm k_{2}\ell_{3}\pm k_{3}\ell_{2}$, $dy_{2}=\pm k_{3}\ell_{1}\pm k_{1}\ell_{3}$, and $dy_{3}=\pm k_{1}\ell_{2}\pm k_{2}\ell_{1}$ holds simultaneously with appropriate choices of signs.
Now we shall state our result in more detail.
Theorem 1.1.
Assume that $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}$ and $r$ are nonzero integers with $x_{1},x_{2},x_{3}>1$, $\{x_{1},x_{2},x_{3}\}\neq\{2,3,7\}$ satisfying (13), and $\gcd(y_{1},y_{2},y_{3})=1$ and $m_{1},m_{2},s_{i},k_{i},\ell_{i}(i=1,2,3)$ are corresponding integers with $m_{2}>m_{1}>0$ satisfying (15).
  • I.
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ divides $x_{i}^{2}+1$ for $i=1,2,3$, then $m_{1}<m_{2}<1.342\times 10^{34},x_{i}<\exp(8380114833)$ and $\left\lvert y_{i}\right\rvert<9.152\times 10^{16}$.
  • II.
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ does not divide $x_{i}^{2}+1$ for some $i$, then $m_{1}<1.017\times 10^{24}$, $m_{2}<\exp(191323)$, $x_{i}<\exp(3.217\times 10^{12})$, and $\left\lvert y_{i}\right\rvert<8.334\times 10^{18}$. Moreover, $m_{2}<e^{40000}$ if additionally $m_{1}\geq 3.537\times 10^{22}$.
  • III.
    If $m_{1}m_{2}<10^{10}$, then $x_{i}<\exp(8813999998)$, and $\left\lvert y_{i}\right\rvert<4.508\times 10^{18}$.
More detailed estimates are given in Tables 5 and 6.
We use a lower bound for linear forms in three logarithms in order to obtain upper bounds for exponents $k_{i}$’s and $\ell_{i}$’s in terms of $m_{1},m_{2}$.
These upper bounds themselves do not give finiteness of $m_{1}$ and $m_{2}$. However, noting that $r\neq 0$, which gives $\left\lvert\sum_{i}y_{i}\arctan(1/x_{i})\right\rvert\geq\pi/4$, the first case can be easily settled using these upper bounds. In order to settle the second case, we additionally need an upper bound $m_{2}$ in terms of $m_{1}$. This can be done using a lower bound for a quantity of the form $y\arctan(1/x)-r\pi/2$, which gives a linear form of two logarithms.
Our PAGI-GP script is available from
https://drive.google.com/drive/folders/1gFTjX05ts5eaz8ncOjfaxRMvCW-drvGX.
(12) can be seen as a special case of the generalized Ramanujan-Nagell equation
(16) $x^{2}+Ax+B=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{n}^{e_{n}},$
where $A$ and $B$ are given integers with $A^{2}-4B\neq 0$ and $p_{1},p_{2},\ldots,p_{n}$ are given primes. Evertse [3] proved that (16) has at most $3\times 7^{4n+6}$ solutions. In the case $n=2$, the author [20] reduced Evertse’s bound $3\times 7^{14}$ to $63$.
On the other hand, our result does not give an upper bound for numbers of solutions
(17) $x^{2}+1=2^{s}p_{1}^{k}p_{2}^{\ell}$
since the case $r=0$ is not considered. Indeed, Størmer [13] implicitly pointed out that, if $x^{2}+1=ay$, then
(18) $\arctan\frac{1}{az-x}-\arctan\frac{1}{az+a-x}=\arctan\frac{1}{az(z+1)-(2z+1)x+y}.$
Størmer [14] showed that (17) has at most one solution with each fixed combination of parities of $s_{i},k_{i},\ell_{i}$ with zero and nonzero-even distinguished. Although there exist $18$ combinations (01,012,012)(0$\mid$ 1,0$\mid$ 1$\mid$ 2,0$\mid$ 1$\mid$ 2)( 0 ∣ 1 , 0 ∣ 1 ∣ 2 , 0 ∣ 1 ∣ 2 ), all-even combinations can clearly be excluded and therefore (17) has at most $14$ solutions totally.
2. Preliminaries
In this section, we introduce some notation and some basic facts.
For any Gaussian integer $\eta$, we have an associate $\eta^{\prime}$ of $\eta$ such that $-\pi/4<\arg\eta^{\prime}<\pi/4$ and therefore $-\pi/2<\arg\eta^{\prime}/\bar{\eta^{\prime}}<\pi/2$. As usual, the absolute logarithmic height $\operatorname*{h}(\alpha)$ of an algebraic number $\alpha$ of degree $d$ over ${\mathbb{Q}}$ is defined by
(19) $\operatorname*{h}(\alpha)=\frac{1}{d}\left(\log\left\lvert a\right\rvert+\sum_{i=1}^{d}\max\{0,\log\left\lvert\alpha^{(i)}\right\rvert\}\right),$
where $a$ is the smallest integer such that $a\alpha$ is an algebraic integer and the $\alpha^{(i)}$’s $(i=1,2,\ldots,d)$ are the complex conjugates of $\alpha$ over ${\mathbb{Q}}$.
We call a formula (7) to be degenerate if
(20) $\sum_{i\in S}y_{i}^{\prime}\arctan\frac{1}{x_{i}}=\frac{r^{\prime}\pi}{4}$
for some proper subset $S$ of $\{1,2,\ldots,n\}$ and integers $y_{i}^{\prime}\leavevmode\nobreak\ (i\in S)$ and $r^{\prime}$ which may be zero but not all zero.
We shall show that a degenerate case of a three-term Machin-type formula occurs only when $\{x_{1},x_{2},x_{3}\}=\{2,3,7\}$.
Lemma 2.1.
Let $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}$ and $r$ be nonzero integers with $x_{1},x_{2},x_{3}>1$ satisfying (13) and $m_{1},m_{2},s_{i},k_{i},\ell_{i}(i=1,2,3)$ be corresponding integers satisfying (15). If $\ell_{1}=\ell_{2}=0$, then $\{x_{1},x_{2},x_{3}\}=\{2,3,7\}$.
Proof.
Let $\eta_{1},\eta_{2}$ be corresponding Gaussian integers in (14). Since $\ell_{1}=\ell_{2}=0$, we have $[(x_{i}+{\sqrt{-1}})/(x_{i}-{\sqrt{-1}})]=[\eta_{1}/\bar{\eta}_{1}]^{\pm k_{i}}$ for $i=1,2$. Hence, we obtain
(21) $k_{2}\arctan\frac{1}{x_{1}}\pm k_{1}\arctan\frac{1}{x_{2}}=\frac{r^{\prime}\pi}{4}$
for some integer $r^{\prime}$, which may be zero. Substituting this into (13) multiplied by an appropriate factor, we have
(22) $y_{1}^{\prime}\arctan\frac{1}{x_{1}}+y_{3}^{\prime}\arctan\frac{1}{x_{3}}=\frac{r^{\prime\prime}\pi}{4}$
for some integers $y_{1}^{\prime},y_{3}^{\prime},r^{\prime\prime}$, which may be zero again but not all zero. Now Størmer’s result in [12] on two-term Machin-type formulae implies that $\{x_{1},x_{2},x_{3}\}=\{2,3,7\}$. ∎
Moreover, we can determine the only solution for (14) with $k_{i}=0$ and $\ell_{i}>2$.
Lemma 2.2.
The equation $x^{2}+1=2^{e}y^{n}$ with $x>0,n>2$ has only one integral solution $(x,e,y,n)=(239,1,13,4)$.
Proof.
From the result of [7], $x^{2}+1=m^{t}$ with $x>0,t>1$ has no solution. Théorème 8 of [14] shows that $x^{2}+1=2m^{t}$, then $t$ must be a power of two. By Ljunggren’s result [8], the only integer solution of $x^{2}+1=2m^{4}$ with $x,m>1$ is $(x,m)=(239,13)$. Easier proofs of Ljunggren’s result have been obtained by Steiner and Tzanakis [11] and Wolfskill [17]. ∎
3. A lower bound for linear forms of three logarithms
Our argument depends on a lower bound for linear forms of three logarithms. Results in a paper [10] entitled a kit are technical but still worthful to use for the purpose of improving our upper bounds. We use Theorem 4.1 of an earlier version [9] rather than [10], which enables us to obtain better bounds. However, the proof of Theorem 4.1 in [9] needs an amendment.
Lemma 3.1.
Let $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ be three distinct algebraic numbers not equal to zero or one, which are either all $>1$ or all complex of modulus $1$. Let $b_{1}$, $b_{2}$, and $b_{3}$ be three positive rational integers with $\gcd(b_{1},b_{2},b_{3})=1$, and put
$\Lambda=b_{1}\log\alpha_{1}+b_{2}\log\alpha_{2}-b_{3}\log\alpha_{3},$
where the logarithms of the $\alpha_{i}$’s are real if $\alpha_{i}>1$ and purely imaginary if $\left\lvert\alpha_{i}\right\rvert=1$. Let $w$ is the maximal order of a root of unity which belongs to the field ${\mathbb{Q}}(\alpha_{1},\alpha_{2},\alpha_{3})$ and assume that $0<\left\lvert\Lambda\right\rvert<2\pi/w$.
We put $d_{i}=\gcd(b_{i},b_{3})$ and $b_{i}=d_{i}b_{i}^{\prime}$ for $i=1,2$ and $b_{3}=d_{1}b_{3}^{\prime}=d_{2}b_{3}^{\prime\prime}$. Let $K\geq 3$, $L\geq 5$, and $R_{i}$ and $R_{ij}$ with $1\leq i,j\leq 3$ be positive rational integers satisfying
$R_{i}>R_{i1}+R_{i2}+R_{i3}$
for $i=1,2,3$. Let $\rho\geq 2$ be a real number and suppose that
(23) $\left(KL-K-\frac{K}{3L}\right)\log\rho\geq(D+1)\log N+gL(a_{1}R_{1}+a_{2}R_{2}+a_{3}R_{3})+D(K-1)\log b,$
where $N=K^{2}L$, $D=[{\mathbb{Q}}(\alpha_{1},\alpha_{2},\alpha_{3})\colon{\mathbb{Q}}]/[{\mathbb{R}}(\alpha_{1},\alpha_{2},\alpha_{3})\colon{\mathbb{R}}]$,
$g=\frac{1}{4}-\frac{N}{12R_{1}R_{2}R_{3}},\leavevmode\nobreak\ b=(b_{3}^{\prime}\eta_{0})(b_{3}^{\prime\prime}\zeta_{0})\left(\prod_{k=1}^{K-1}k!\right)^{-4/(K^{2}-K)}$
with
$\eta_{0}=\frac{R_{1}-1}{2}+\frac{(R_{3}-1)b_{1}}{2b_{3}},\leavevmode\nobreak\ \zeta_{0}=\frac{R_{2}-1}{2}+\frac{(R_{3}-1)b_{2}}{2b_{3}},$
and
$a_{i}\geq\rho\left\lvert\log\alpha_{i}\right\rvert-\log\left\lvert\alpha_{i}\right\rvert+2D\operatorname*{h}(\alpha_{i})$
for $i=1,2,3$. Put
$V=\sqrt{(R_{11}+1)(R_{21}+1)(R_{31}+1)}.$
If, for some positive real number $\chi$,
  • (i)
    $(R_{11}+1)(R_{21}+1)(R_{31}+1)>KM$, where
    $M=\max\{R_{11}+R_{21}+1,R_{21}+R_{31}+1,R_{31}+R_{11}+1,\chi V\},$
  • (ii)
    #$\#\{\prod_{i=1}^{3}\alpha_{i}^{r_{i}}:0\leq r_{i}\leq R_{i1}\}>L$,
  • (iii)
    $(R_{12}+1)(R_{22}+1)(R_{32}+1)>2K^{2}$,
  • (iv)
    #$\#\{\prod_{i=1}^{3}\alpha_{i}^{r_{i}}:0\leq r_{i}\leq R_{i2}\}>2KL$, and
  • (v)
    $(R_{13}+1)(R_{23}+1)(R_{33}+1)>6K^{2}L$,
then either
  • (A)
    $\Lambda^{\prime}>\rho^{-KL}$, where
    $\Lambda^{\prime}=\left\lvert\Lambda\right\rvert\frac{LR_{3}e^{LR_{3}\left\lvert\Lambda\right\rvert/(2b_{3})}}{2b_{3}},$
  • (B)
    $b_{i}\leq\max\{R_{i1},R_{i2}\}$ for each $i=1,2,3$, or
  • (C)
    there exist rational integers $u_{i}$, not all zero, such that
    $u_{1}b_{1}+u_{2}b_{2}+u_{3}b_{3}=0$
    with $\gcd(u_{1},u_{2},u_{3})=1$ and
    $\left\lvert u_{i}\right\rvert\leq\frac{(R_{i+1,1}+1)(R_{i+2,1}+1)}{M-\max\{R_{i+1,1},R_{i+2,1}\}}$
    for each $i=1,2,3$, where the indices are reduced modulo $3$.
Proof.
We use notation such as $p_{i}(z_{1},z_{2})$, $\Psi_{\mathcal{I}}(x)$, and $\Theta(D,{\mathcal{I}})$ according to [9]. In p.8, l.-7 of [9], they wrote that for a certain linear transformation $T$, the identity
$T(\theta_{1}z_{1}+\theta_{2}z_{2})=\theta_{1}(z_{1}+z_{2})+\theta_{2}(-\theta_{1}z_{1}/\theta_{2})=\theta_{1}z_{1}$
holds, where the right hand side $\theta_{1}z_{1}$ should be $\theta_{1}z_{2}$. Thus, $p_{i}(z_{1},z_{2})$ cannot be reduced to a polynomial $p_{i}^{(T)}(z_{1},z_{2})$ with degree at most $D_{1}$ in $z_{2}$ in their way.
However, the degree of $p_{i}^{(T)}(z_{1},z_{2})$ is at most $\max\{D_{1},D_{2}\}$ and, Lemma 3.5 in [9] holds with $\Theta(\max\{D_{1},D_{2}\},\#{\mathcal{I}})$ in place of $\max\{\Theta(D_{1},\#{\mathcal{I}}),\Theta(D_{2},\#{\mathcal{I}})\}$. This suffices for our purpose since we set $D_{1}=D_{2}=K-1$.
More explicitly, if $\theta_{1}\theta_{2}\neq 0$, then we set $\eta_{j}^{\prime}=\eta_{j}+\theta_{2}\zeta_{j}/\theta_{1}$ and $\zeta_{j}^{\prime}=-\theta_{2}\eta_{j}/\theta_{1}$, and $q_{i}(z_{1},z_{2})=p_{i}(z_{1}+z_{2},-\theta_{1}z_{2}/\theta_{2})$. Then we have $q_{i}$ are polynomials with degree $\max\{D_{1},D_{2}\}$ in $z_{2}$ such that $q_{i}(x\eta_{j}^{\prime},x\zeta_{j}^{\prime})=p_{i}(x\eta_{j},x\zeta_{j})$ for $i=1,\ldots,N$. Hence, for $1\leq i,j\leq N$, we have
(24) $p_{i}(x\eta_{j},x\zeta_{j})f_{i}(x(\theta_{1}\eta_{j}+\theta_{2}\zeta_{j})=q_{i}(x\eta_{j}^{\prime},x\zeta_{j}^{\prime})f_{i}(\theta x\eta_{j}^{\prime})$
with $\theta=\theta_{1}$ and each $q_{i}$ has degree at most $\max\{D_{1},D_{2}\}$ in $z_{2}$.
It may be better to give an explicit argument in the case $\theta_{1}\theta_{2}=0$. If $\theta_{1}=0$, then, for each $1\leq i,j\leq N$, putting $\eta_{j}^{\prime}=\zeta_{j}$, $\zeta_{j}^{\prime}=\eta_{j}$, and $q_{i}(z_{1},z_{2})=p_{i}(z_{2},z_{1})$, we have (24) with $\theta=\theta_{2}$ and each $q_{i}$ has degree at most $D_{1}$ in $z_{2}$. Similarly, if $\theta_{2}=0$, then, for each $1\leq i,j\leq N$, putting $\eta_{j}^{\prime}=\eta_{j}$, $\zeta_{j}^{\prime}=\zeta_{j}$, and $q_{i}(z_{1},z_{2})=p_{i}(z_{1},z_{2})$, we have (24) with $\theta=\theta_{1}$ and each $q_{i}$ has degree at most $D_{2}$ in $z_{2}$.
Now, (24) yields that
(25) $\Psi_{\mathcal{I}}(x)=\det(\delta_{i,j}q_{i}(x\eta_{j}^{\prime},x\zeta_{j}^{\prime})f_{i}(\theta x\eta_{j}^{\prime})$
and $\Psi_{\mathcal{I}}(x)$ has a zero at $x=0$ of multiplicity at least $\Theta(\max\{D_{1},D_{2}\},\#{\mathcal{I}})$.
For the remaining part of the proof, we can proceed as in [10] with changes of constants according to [9] and using Laurent’s zero lemma, which is given as Proposition A.1 in [10] in the place of Proposition 3.11 in [10]. ∎
We also need a lower bound for linear forms in two logarithms.
Lemma 3.2.
Let $h,\rho$ and $\mu$ be real numbers with $\rho>1,1/3\leq\mu\leq 1$. Set
$\sigma=\frac{1+2\mu-\mu^{2}}{2},\lambda=\sigma\log\rho,H=\frac{h}{\lambda}+\frac{1}{\sigma},\omega=2\left(1+\sqrt{1+\frac{1}{4H^{2}}}\right),\theta=\frac{1}{2H}+\sqrt{1+\frac{1}{4H^{2}}}.$
Let
(26) $\Lambda_{1}=b_{1}\log\alpha_{1}+b_{2}\log\alpha_{2},$
where $b_{1},b_{2}$ are positive integers and $\alpha_{1}=(X_{1}+Y_{1}{\sqrt{-1}})/(X_{1}-Y_{1}{\sqrt{-1}}),\alpha_{2}=(X_{2}+Y_{2}{\sqrt{-1}})/(X_{2}-Y_{2}{\sqrt{-1}})$ are Gaussian rationals with $X_{i}>\left\lvert Y_{i}\right\rvert>0$ integers. Put $N_{i}=X_{i}^{2}+Y_{i}^{2},a_{i}^{\prime}=\rho_{1}\left\lvert\log\alpha_{i}\right\rvert+2\operatorname*{h}(\alpha_{i})=\rho_{1}\left\lvert\log\alpha_{i}\right\rvert+\log N_{i}$ (we use this notation here in order to avoid confusion with $a_{i}$) and
(27) ${b^{\prime\prime}}=\frac{\left\lvert b_{1}\right\rvert}{a_{2}^{\prime}}+\frac{\left\lvert b_{2}\right\rvert}{a_{1}^{\prime}}.$
Furthermore, assume that $a_{1}^{\prime}a_{2}^{\prime}\geq\lambda^{2}$ and $h\geq\log{b^{\prime\prime}}+\log\lambda+1.81$.
Then we have
(28) $\log\left\lvert\Lambda_{1}\right\rvert>-c\left(h+\frac{\lambda}{\sigma}\right)^{2}a_{1}^{\prime}a_{2}^{\prime}-\sqrt{\omega\theta}\left(h+\frac{\lambda}{\sigma}\right)-\log\left(c^{\prime}\left(h+\frac{\lambda}{\sigma}\right)^{2}a_{1}^{\prime}a_{2}^{\prime}\right),$
where
(29) $c=\frac{\mu}{\lambda^{3}\sigma}\left(\frac{\omega}{6}+\sqrt{\frac{\omega^{2}}{36}+\frac{2\lambda\omega^{5/4}\theta^{1/4}}{3\sqrt{a_{1}^{\prime}a_{2}^{\prime}H}}+\frac{\lambda\omega}{3H}\left(\frac{1}{a_{1}^{\prime}}+\frac{1}{a_{2}^{\prime}}\right)}\right)^{2},c^{\prime}=\sqrt{\frac{c\sigma\omega\theta}{\lambda^{3}\mu}}.$
Proof.
We apply Theorem 2 of [6] with $D=1$ and $h=\log{b^{\prime\prime}}+\log\lambda+1.81$. We note that our settings always yield that $\log{b^{\prime\prime}}+\log\lambda+1.81>\max\{\lambda,(\log 2)/2\}$ in Theorem 2 of [6]. Thus, we obtain (28). ∎
Now we are to describe our explicit bounds.
Theorem 3.3.
Let $B_{0}=e^{20}$ and $X_{1},Y_{1},X_{2},Y_{2}$ be nonzero rational integers such that $m_{i}=X_{i}^{2}+Y_{i}^{2}$ are odd, $\gcd(X_{i},Y_{i})=1$ for both $i=1,2$, and neither of $m_{1}$ nor $m_{2}$ is a perfect power. Moreover assume that $m_{1}\leq 41$ and $m_{2}\geq\tilde{m}_{2}$ with $\tilde{m}_{2}$ from Table 1 or $m_{i}\geq\tilde{m}_{i}$ for $i=1,2$ with $\tilde{m}_{1}$ and $\tilde{m}_{2}$ from Table 1. Write $\eta_{i}=(X_{i}+Y_{i}{\sqrt{-1}})/(X_{i}-Y_{i}{\sqrt{-1}})$ for $i=1,2$ and $\eta_{3}={\sqrt{-1}}$ and put
(30) $\Lambda=\pm e_{1}\log\eta_{1}\pm e_{2}\log\eta_{2}\pm e_{3}\log\eta_{3}.$
Writing $\Omega=(\log m_{1})(\log m_{2})$ and $E=(e_{1}\log m_{1}+e_{2}\log m_{2})/\Omega$, we have either $E<A_{0}$, $(A_{2}E)^{2}\log m_{1}<B_{0}$,
(31) $-\log\left\lvert\Lambda\right\rvert<A_{1}\log^{2}(A_{2}E\sqrt{\log m_{1}})\Omega,$
or
(32) $e_{1}\log m_{1}+e_{2}\log m_{2}<A_{3}\Omega\log(A_{2}E\sqrt{\log m_{1}}),$
where $A_{1},A_{2},A_{3}$, and $\rho$ are constants given in Table 1. Here and hereafter, we write $m_{1}^{*}=\max\{m_{1},10^{10}/m_{1}\}$.
Table 1. Constants in bounds for Theorem 3.3
$m_{1}$ $\tilde{m}_{2}$ $A_{1}$ $A_{2}$ $A_{3}$ $A_{0}$ $\rho$
$5$ $2000000005$ $6792.639$ $1.987228$ $356.1620$ $113.2663$ $4.531$
$13$ $769230773$ $5831.233$ $1.933988$ $265.7853$ $84.8351$ $4.635$
$17$ $588235297$ $3300.665$ $1.735844$ $275.7772$ $75.8539$ $5.155$
$29$ $344827589$ $3812.426$ $1.820629$ $242.0317$ $71.5043$ $4.931$
$37$ $270270277$ $2561.555$ $1.583766$ $258.3560$ $63.9328$ $5.661$
$41$ $243902441$ $5072.693$ $1.890607$ $226.4842$ $69.5800$ $4.770$
$5$ $10^{16}$ $6067.262$ $1.757445$ $359.7087$ $107.0678$ $4.791$
$13$ $10^{16}$ $5143.961$ $1.702538$ $274.8311$ $78.7827$ $4.932$
$17$ $10^{16}$ $2865.651$ $1.500984$ $290.9531$ $69.8349$ $5.572$
$29$ $10^{16}$ $3295.501$ $1.566799$ $252.8907$ $65.0122$ $5.336$
$37$ $10^{16}$ $2176.447$ $1.410974$ $258.1061$ $59.7321$ $5.878$
$41$ $10^{16}$ $4394.894$ $1.620517$ $235.0623$ $62.7626$ $5.169$
$5$ $10^{24}$ $5719.809$ $1.703384$ $356.2457$ $105.9955$ $4.799$
$13$ $10^{24}$ $4843.038$ $1.649921$ $271.1685$ $77.7108$ $4.938$
$17$ $10^{24}$ $2679.811$ $1.418838$ $283.1656$ $68.1770$ $5.660$
$29$ $10^{24}$ $3089.378$ $1.482171$ $245.5286$ $63.3267$ $5.421$
$37$ $10^{24}$ $2028.535$ $1.331976$ $250.9353$ $58.1814$ $5.975$
$41$ $10^{24}$ $4129.103$ $1.569853$ $231.1519$ $61.7086$ $5.172$
$5$ $e^{40000}$ $4994.390$ $1.479858$ $375.2430$ $100.3593$ $5.135$
$13$ $e^{40000}$ $4214.314$ $1.431915$ $271.9920$ $72.8793$ $5.249$
$17$ $e^{40000}$ $2287.357$ $1.189845$ $307.5275$ $62.8590$ $6.243$
$29$ $e^{40000}$ $2652.249$ $1.242648$ $264.3578$ $57.7106$ $5.986$
$37$ $e^{40000}$ $1701.715$ $1.069216$ $297.9336$ $51.8596$ $6.940$
$41$ $e^{40000}$ $3591.493$ $1.289024$ $252.7466$ $55.1890$ $5.834$
$\tilde{m}_{1}$ $\tilde{m}_{2}$ $A_{1}$ $A_{2}$ $A_{3}$ $A_{0}$ $\rho$
$53$addstackgap[.5]$m_{1}^{*}$ $5381.263$ $1.985977$ $219.4190$ $69.6043$ $4.589$
$53$ $10^{24}$ $4370.434$ $1.638498$ $214.9439$ $61.5175$ $4.965$
$53$ $10^{32}$ $4230.482$ $1.546752$ $221.6378$ $59.1862$ $5.166$
$53$ $10^{34}$ $4205.618$ $1.541857$ $221.2940$ $59.0913$ $5.167$
$53$ $e^{2000}$ $3811.172$ $1.370913$ $221.6077$ $55.4326$ $5.473$
$53$ $e^{40000}$ $3794.956$ $1.398933$ $224.5476$ $55.9594$ $5.397$
$10^{24}$ $10^{24}$ $1469.356$ $1.105031$ $139.9162$ $24.0442$ $7.292$
$5000$ $e^{2000}$ $2375.090$ $1.189591$ $177.7400$ $36.3032$ $6.256$
$5000$ $e^{40000}$ $2363.495$ $1.186121$ $177.4782$ $36.2441$ $6.259$
$10^{7}$ $e^{40000}$ $1742.083$ $1.057810$ $154.1620$ $27.6471$ $6.957$
$10^{14}$ $e^{40000}$ $1368.425$ $0.943348$ $146.8828$ $21.8950$ $7.836$
Proof.
We may assume that $X_{i}>Y_{i}>0$ for $i=1,2$ and $\left\lvert\Lambda\right\rvert<10^{-9}$. We can take $\alpha_{1}=\pm{\sqrt{-1}}$, $\alpha_{2}=\eta_{1}^{\pm 1}$, and $\alpha_{3}=\eta_{2}^{\pm 1}$ with a suitable choice of the combination of signs such that
(33) ±$\pm\Lambda=e_{3}\log\alpha_{1}+e_{1}\log\alpha_{2}-e_{2}\log\alpha_{3}$
with $(b_{1},b_{2},b_{3})=(e_{3},e_{1},e_{2})$.
Now we set $\rho$ according to Table 1 and take $a_{1}=\rho\pi/2$ and $a_{3}=\rho\pi/2+\log m_{2}$. Moreover, we can take $a_{2}=2\rho\arctan(Y_{1}/X_{1})+\log m_{1}$ if $m_{1}\leq 41$ and $a_{2}=\rho\pi/2+\log m_{1}$. Hence, we have $\rho>4$, $a_{1}\geq 2\pi$, $a_{2}\geq 6\arctan(1/4)+\log 17=4.79304\cdots$. Moreover, we see that $a_{2}a_{3}\geq(9.062\arctan(1/4)+\log 37)(2.2655\pi+\log 270270277)=135.5131\cdots$. We note that in the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*}\})$, we have $a_{2}a_{3}\geq(\log(53)+\rho\pi/2)(\log(10^{10}/53)+\rho\pi/2)$ since $(\log x+\rho\pi/2)(\log(10^{10})-\log x+\rho\pi/2)$ is increasing in the range $1\leq x\leq 10^{5}$.
We set $\chi$, $C_{0}$, and $C_{1}$ according to Table 2, $C_{2}=6^{1/3}$ and put
(34) $c_{1}^{\prime}=(\chi_{1}C_{1})^{2/3}+\frac{C_{2}C_{0}^{1/3}C_{1}^{2/3}}{\log^{1/3}\rho}+\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}}$
and
(35) $c_{i}^{\prime}=(\chi_{1}C_{1})^{2/3}+\frac{C_{2}C_{0}^{1/3}C_{1}^{2/3}}{\log^{1/3}\rho}+\max\left\{\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}},\sqrt{\frac{C_{0}C_{1}}{a_{5-i}^{2}a_{1}\log\rho}}\right\}$
for $i=2,3$. Moreover, we set
(36) ${b^{\prime}}=\left(\frac{c_{1}b_{3}^{\prime}}{a_{1}}+\frac{c_{3}b_{1}^{\prime}}{a_{3}}\right)\left(\frac{c_{2}b_{3}^{\prime\prime}}{a_{2}}+\frac{c_{3}b_{2}^{\prime}}{a_{3}}\right)$
and $b_{0}=(A_{2}E)^{2}\log m_{1}$. Finally, we put $K=\left\lfloor C_{1}a_{1}a_{2}a_{3}\log b_{0}\right\rfloor$, $L=\left\lfloor C_{0}\log b_{0}/\log\rho\right\rfloor$, and $\chi=\chi_{1}\log^{1/2}b_{0}$. We note that $K\geq K_{0}$ and $L\geq L_{0}$ for constants $K_{0}$ and $L_{0}$ in Table 2 in each case.
We write $i=k+1,j=k+2$ for brevity and put $R_{i1}=\left\lfloor(\chi^{2}K^{2}a_{i+1}a_{i+2}/a_{i}^{2})^{1/3}\right\rfloor$ for $i=1,2,3$.
Since $\chi_{1}>0.6$, $C_{1}>6$, $K>540000$, $C_{0}\leq 3.5$, and $a_{1}a_{2}a_{3}>20$ by our settings and $b_{0}\geq B_{0}=e^{20}$ by assumption, we have $((1-1/K)\chi_{1}C_{1})^{2/3}>C_{0}^{1/2}/(2\sqrt{a_{1}a_{2}a_{3}\log^{1/2}b_{0}})$ and therefore
(37) $R_{k+1,1}+1>(\chi^{2}(K-1)^{2}a_{k}a_{k+2}/a_{k+1}^{2})^{1/3}>\sqrt{\frac{C_{0}a_{k+2}\log b_{0}}{4a_{k+1}}}$
and, similarly,
(38) $R_{k+2,1}+1>\sqrt{\frac{C_{0}a_{k+1}\log b_{0}}{4a_{k+2}}}.$
Hence, we obtain $4(R_{k+1,1}+1)(R_{k+2,1}+1)>L$ to confirm (ii).
Observing that $\chi_{1}>0.6$, $C_{1}>6$, $K>540000$, $C_{0}\leq 3.5$, $a_{i}>2\pi$ for $i=1,2,3$, and $b_{0}\geq B_{0}=e^{20}$, we have
(39) $R_{i1}+R_{i+1,1}+1<1+(\chi_{1}C_{1})^{2/3}a_{i+2}(a_{i}+a_{i+1})\log b_{0}<\chi_{1}C_{1}a_{1}a_{2}a_{3}\log b_{0}$
for $i=1,2,3$ and therefore
(40) $\begin{split}(R_{11}+1)(R_{21}+1)(R_{31}+1)>&\leavevmode\nobreak\ (\chi_{1}C_{1}a_{1}a_{2}a_{3})^{2}\log^{3}b_{0}>K\chi_{1}^{2}C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0}\\ >&\leavevmode\nobreak\ K\max\{R_{11}+R_{21}+1,R_{21}+R_{31}+1,R_{31}+R_{11}+1\}.\end{split}$
Observing that $(R_{11}+1)(R_{21}+1)(R_{31}+1)>(\chi K)^{2}$, we obtain $(R_{11}+1)(R_{21}+1)(R_{31}+1)>\chi KV$. Hence, we can confirm (i).
Nextly, we put
$\begin{split}R_{12}=&\leavevmode\nobreak\ \left\lfloor(2K^{2}a_{2}a_{3}/a_{1}^{2})^{1/3}\right\rfloor,\\ R_{22}=&\leavevmode\nobreak\ \left\lfloor\max\left\{\sqrt{\frac{KLa_{3}}{2a_{2}}},\left(\frac{2K^{2}a_{1}a_{3}}{a_{2}^{2}}\right)^{1/3}\right\}\right\rfloor,R_{32}=\left\lfloor\max\left\{\sqrt{\frac{KLa_{2}}{2a_{3}}},\left(\frac{2K^{2}a_{1}a_{2}}{a_{3}^{2}}\right)^{1/3}\right\}\right\rfloor.\end{split}$
We immediately see that $4(R_{22}+1)(R_{32}+1)>2KL$ and $(R_{12}+1)(R_{22}+1)(R_{32}+1)>2K^{2}$ to confirm (iii) and (iv) respectively.
Finally, we put $R_{i3}=\left\lfloor C_{2}(K^{2}La_{i+1}a_{i+2}/a_{i}^{2})^{1/3}\right\rfloor$ for $i=1,2,3$. We immediately have (v) since $C_{2}\geq 6^{1/3}$.
In each case, noting that $a_{k+2}>a_{k+1}>a_{k}$, we have $R_{k,j}\geq R_{k+1,j}\geq R_{k+2,j}$ for $j=1,3$ and $R_{k+2,j}$ can be minorized by replacing $m_{2}$ by $\tilde{m}_{2}$ and then $m_{1}$ by $\tilde{m}_{1}$. In the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$, $R_{k+2,j}$ can be minorized by replacing $m_{2}$ by $10^{10}/m_{1}$ and then taking $m_{1}=53$. Moreover, we observe that $R_{i2}\geq\left\lfloor(2K^{2}a_{k}a_{i}/a_{j}^{2})^{1/3}\right\rfloor$, which can be minorized by the same way. Hence, we have $R_{1}+R_{2}+R_{3}\geq R_{0}$ for $i=1,2,3$ for constants $K_{0}$ and $L_{0}$ in Table 2 in each case.
Now we have
(41) $\sum_{i=1}^{3}a_{i}R_{i1}<3(\chi_{1}C_{1})^{2/3}a_{1}a_{2}a_{3}\log b_{0},$
either
(42) $\sum_{i=1}^{3}a_{i}R_{i2}<a_{1}a_{2}a_{3}\left(2\sqrt{\frac{2C_{0}C_{1}}{\rho\pi\log\rho}}\log b_{0}+2^{1/3}C_{1}^{2/3}\log^{2/3}b_{0}\right)$
or
(43) $\sum_{i=1}^{3}a_{i}R_{i2}<3\times 2^{1/3}C_{1}^{2/3}a_{1}a_{2}a_{3}\log^{2/3}b_{0},$
and
(44) $\sum_{i=1}^{3}a_{i}R_{i3}<3C_{2}C_{0}^{1/3}C_{1}^{2/3}a_{1}a_{2}a_{3}\log b_{0}/\log^{1/3}\rho.$
Taking $R_{i}=R_{i1}+R_{i2}+R_{i3}+1$ and writing $c_{4}=(1+1/R_{0})^{3}$, we see that
(45) $\begin{split}&R_{1}R_{2}R_{3}=\prod_{i=1}^{3}(R_{i1}+R_{i2}+R_{i3}+1)\\ <&\leavevmode\nobreak\ c_{4}((\chi K)^{2/3}+C_{2}K^{2/3}L^{1/3}+\sqrt{KL}a_{1}^{-1/3}a_{2}^{1/6}a_{3}^{1/6})^{2}\\ &((\chi K)^{2/3}+C_{2}K^{2/3}L^{1/3}+(2K^{2})^{1/3})\\ <&\leavevmode\nobreak\ c_{4}\left(C_{2}+\frac{\chi_{1}^{2/3}\log^{1/3}b_{0}}{L^{1/3}}+\left(\frac{L}{K}\right)^{1/6}a_{1}^{-1/3}a_{2}^{1/6}a_{3}^{1/6}\right)^{2}\\ &\left(C_{2}+\frac{2^{1/3}+\chi_{1}^{2/3}\log^{1/3}b_{0}}{L^{1/3}}\right)K^{2}L\\ <&\leavevmode\nobreak\ c_{4}\left(C_{2}+\left(\frac{L\log\rho}{(L-1)C_{0}}\right)^{1/3}\chi_{1}^{2/3}+\left(\frac{KC_{0}}{C_{1}(K-1)\log\rho}\right)^{1/6}a_{1}^{-1/2}\right)^{2}\\ &\left(C_{2}+\left(\frac{L\log\rho}{(L-1)C_{0}}\right)^{1/3}\chi_{1}^{2/3}+\left(\frac{2}{L}\right)^{1/3}\right)K^{2}L\end{split}$
or
(46) $\begin{split}R_{1}R_{2}R_{3}<c_{4}^{3}\left(C_{2}+\left(\frac{L\log\rho}{(L-1)C_{0}}\right)^{1/3}\chi_{1}^{2/3}+\left(\frac{2}{L}\right)^{1/3}\right)^{3}K^{2}L.\end{split}$
Writing $R_{1}R_{2}R_{3}<c_{0}K^{2}L$, we have $g<1/4-1/(12c_{0})\leq g_{0}$ with $g_{0}$ taken from Table 3.
Thus, we conclude that
(47) $gL\sum_{i=1}^{3}a_{i}R_{i}<g_{0}L(a_{1}+a_{2}+a_{3}+C_{3}^{\prime}a_{1}a_{2}a_{3}\log b_{0})<C_{3}a_{1}a_{2}a_{3}\log^{2}b_{0},$
where
(48) $C_{3}^{\prime}=3(\chi_{1}C_{1})^{2/3}+\frac{3C_{2}C_{0}^{1/3}C_{1}^{2/3}}{\log^{1/3}\rho}+\max\left\{2\sqrt{\frac{2C_{0}C_{1}}{\rho\pi\log\rho}}+\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}},\frac{3\times 2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}}\right\}$
and
(49) $C_{3}=\frac{g_{0}C_{0}}{\log\rho}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}\log^{2}B_{0}}+C_{3}^{\prime}\right).$
Now we would like to show that $\log b\leq\log b_{0}$.
We see that $R_{i}-1=R_{i1}+R_{i2}+R_{i3}<c_{i}^{\prime}a_{i+1}a_{i+2}\log b_{0}$ for $i=1,2,3$. Hence, we obtain
(50) $b_{3}^{\prime}\eta_{0}=\frac{(R_{1}-1)b_{3}^{\prime}}{2}+\frac{(R_{3}-1)b_{1}b_{3}^{\prime}}{2b_{3}}<\frac{\log b_{0}}{2}(c_{1}^{\prime}a_{2}a_{3}b_{3}^{\prime}+c_{3}^{\prime}a_{1}a_{2}b_{1}^{\prime}).$
and
(51) $b_{3}^{\prime\prime}\zeta_{0}=\frac{(R_{2}-1)b_{3}^{\prime\prime}}{2}+\frac{(R_{3}-1)b_{2}b_{3}^{\prime\prime}}{2b_{3}}<\frac{\log b_{0}}{2}(c_{2}^{\prime}a_{3}a_{1}b_{3}^{\prime\prime}+c_{3}^{\prime}a_{1}a_{2}b_{2}^{\prime\prime}).$
Thus, we have
(52) $\begin{split}\log b\leq&\leavevmode\nobreak\ \log\left(\frac{\log^{2}b_{0}}{4}\right)+\log(c_{1}^{\prime}a_{2}a_{3}b_{3}^{\prime}+c_{3}^{\prime}a_{1}a_{2}b_{1}^{\prime})(c_{2}^{\prime}a_{3}a_{1}b_{3}^{\prime\prime}+c_{3}^{\prime}a_{1}a_{2}b_{2}^{\prime\prime})\\ &\leavevmode\nobreak\ -\frac{4}{K(K-1)}\log\left(\prod_{k=1}^{K-1}k!\right).\end{split}$
We see that the last term is greater than
(53) $2\log K-3+\frac{2\log(2\pi K/e^{3/2})}{K-1}-\frac{2+6\pi^{-2}+\log K}{3K(K-1)}>2\log(K+1)-3,$
observing that
(54) $\frac{2+6\pi^{-2}+\log K}{3K(K-1)}-\frac{1+2\log(2\pi K/e^{3/2})}{K-1}<0$
for $K>1$. Hence, we have
(55) $\begin{split}\log b\leq&\log\left(\frac{\log^{2}b_{0}}{4}\right)+\log\left((c_{1}^{\prime}a_{2}a_{3}b_{3}^{\prime}+c_{3}^{\prime}a_{1}a_{2}b_{1}^{\prime})(c_{2}^{\prime}a_{3}a_{1}b_{3}^{\prime\prime}+c_{3}^{\prime}a_{1}a_{2}b_{2}^{\prime\prime})\right)\\ &-2\log(a_{1}a_{2}a_{3}\log b_{0})+3-2\log C_{1}\\ <&\log\left[\left(\frac{c_{1}^{\prime}b_{3}^{\prime}}{a_{1}}+\frac{c_{3}^{\prime}b_{1}^{\prime}}{a_{3}}\right)\left(\frac{c_{2}^{\prime}b_{3}^{\prime\prime}}{a_{2}}+\frac{c_{3}^{\prime}b_{2}^{\prime}}{a_{3}}\right)\right]+3-2\log(2C_{1})<\log{b^{\prime}}.\end{split}$
We see that $c_{i}^{\prime}<A_{2}^{\prime 2}$ for $i=1,2,3$ and therefore ${b^{\prime}}<A_{2}^{\prime 2}{\tilde{b}}$, where $A_{2}^{\prime}$ is taken from Table 3 and
(56) ${\tilde{b}}=\left(\frac{b_{3}}{a_{1}}+\frac{b_{1}}{a_{3}}\right)\left(\frac{b_{3}}{a_{2}}+\frac{b_{2}}{a_{3}}\right).$
Observing that $(b_{2},b_{3})=(e_{1},e_{2})$ and $b_{1}=e_{3}\leq(2/\pi)(e_{1}\theta_{1}+e_{2}\theta_{2})+\left\lvert\Lambda\right\rvert<e_{1}+e_{2}+10^{-9}$, we have
(57) $\begin{split}{\tilde{b}}=&\left(\frac{2e_{2}}{\rho\pi}+\frac{e_{3}}{\log m_{2}}\right)\left(\frac{e_{2}}{\log m_{1}}+\frac{e_{1}}{\log m_{2}}\right)\\ <&\left(\frac{2e_{2}}{\rho\pi}+\frac{e_{1}+e_{2}+10^{-9}}{\log m_{2}}\right)\left(\frac{e_{2}}{\log m_{1}}+\frac{e_{1}}{\log m_{2}}\right)\\ <&(1+10^{-9})\left(\frac{e_{2}}{\log m_{1}}+\frac{e_{1}}{\log m_{2}}\right)\left(e_{2}\left(\frac{2}{\rho\pi}+\frac{1}{\log m_{2}}\right)+\frac{e_{1}}{\log m_{2}}\right)\\ <&\psi^{2}(m_{1},m_{2})\left(\frac{e_{2}}{\log m_{1}}+\frac{e_{1}}{\log m_{2}}\right)^{2},\end{split}$
where
(58) $\psi=\psi(m_{1},m_{2})=\sqrt{(1+10^{-9})\max\left\{\frac{2\log m_{1}}{\rho\pi}+\frac{\log m_{1}}{\log m_{2}},1\right\}}.$
Hence, we have ${\tilde{b}}<\psi^{2}(m_{1},m_{2})E^{2}$ and, combining with (55), we see that ${b^{\prime}}<A_{2}^{\prime 2}{\tilde{b}}<(A_{2}^{\prime}\psi E)^{2}$. Observing that
(59) $(A_{2}^{\prime}\psi)^{2}<A_{2}^{\prime 2}(1+10^{-9})\left(\frac{2}{\rho\pi}+\frac{1}{\log\tilde{m}_{2}}\right)\log m_{1}<A_{2}^{2}\log m_{1},$
we can confirm that ${b^{\prime}}<(A_{2}E)^{2}\log m_{1}=b_{0}$.
Now, observing that $D=1$, it suffices to confirm that
(60) $\left(1-\frac{1}{K}\right)\left(1-\frac{1}{L}-\frac{1}{3L^{2}}\right)C_{0}C_{1}\geq\frac{2\log(C_{0}(C_{1}a_{1}a_{2}a_{3})^{2}\log^{2}b_{0}/\log\rho)}{a_{1}a_{2}a_{3}\log^{2}b_{0}}+C_{1}+C_{3}$
instead of (23). Indeed, we see that the left hand side of (60) is greater than $Y_{0}$ while the right hand side of (60) is smaller than $X_{0}$, where we take $X_{0}$ and $Y_{0}$ from Table 3. Since $X_{0}<Y_{0}$ in any case, we confirm (60). So that, either of (A), (B), or (C) in Lemma 3.1 holds.
In the case (A), we have
(61) $\log\left\lvert\Lambda^{\prime}\right\rvert>-KL\log\rho>-C_{0}C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0}.$
Now we shall show that $-\log\left\lvert\Lambda\right\rvert<(1+10^{-50})KL\log\rho$. We may assume that $-\log\left\lvert\Lambda\right\rvert\geq KL\log\rho$. Since $KL>10^{7}$ and $\rho>3$, we have
(62) $\frac{LR_{3}\left\lvert\Lambda\right\rvert}{2b_{3}}<5KL\times 3^{-KL}<10^{-1000}KL$
and
(63) $\log\frac{LR_{3}}{2b_{3}}<\log\frac{LR_{3}}{2}<10^{-1000}KL.$
Hence, we obtain
(64) $-\log\left\lvert\Lambda\right\rvert<(1+10^{-50})KL\log\rho<(1+10^{-50})C_{0}C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0}<\frac{A_{1}}{4}\Omega\log^{2}b_{0}.$
Recalling that $b_{0}=(A_{2}E)^{2}\log m_{1}$, we have $-\log\left\lvert\Lambda\right\rvert<A_{1}\log^{2}(A_{2}E\sqrt{m_{1}})\Omega$ to obtain (31), where we note that, in the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$,
(65) $\frac{a_{1}a_{2}}{\Omega}=\left(1+\frac{\rho\pi}{2\log m_{1}}\right)\left(1+\frac{\rho\pi}{2\log m_{2}}\right)<\left(1+\frac{3.975\pi}{2\log 53}\right)\left(1+\frac{3.975\pi}{2\log(10^{10}/53)}\right).$
In the case (B), we have
(66) $b_{i}\leq\max\left\{(\chi_{1}C_{1})^{2/3}a_{i+1}a_{i+2}\log b_{0},\sqrt{\frac{C_{0}C_{1}a_{k}}{\log\rho}}a_{j}\log b_{0},(2C_{1}^{2}\log^{2}b_{0})^{1/3}a_{i+1}a_{i+2}\right\}$
for $i\neq j\neq k\neq i$. For $e_{1}=b_{i}$, observing that $a_{i+1}\leq\log m_{2}+\rho\pi/2$ and $a_{i+2}=\rho\pi/2$, we have
(67) $e_{1}\leq\max\left\{\frac{(\chi_{1}C_{1})^{2/3}\rho\pi}{2},\sqrt{\frac{C_{0}C_{1}\rho\pi}{4\log\rho}},\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}}\right\}\left(\log m_{2}+\frac{\rho\pi}{2}\right)\log b_{0}.$
Similarly, we have
(68) $e_{2}\leq\max\left\{\frac{(\chi_{1}C_{1})^{2/3}\rho\pi}{2},\sqrt{\frac{C_{0}C_{1}\rho\pi}{4\log\rho}},\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}}\right\}\left(\log m_{1}+\frac{\rho\pi}{2}\right)\log b_{0}.$
It follows that
(69) $\begin{split}e_{1}\log m_{1}+e_{2}\log m_{2}<&\leavevmode\nobreak\ \max\left\{\frac{(\chi_{1}C_{1})^{2/3}\rho\pi}{2},\sqrt{\frac{C_{0}C_{1}\pi}{4\log\rho}},\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}}\right\}\\ &\leavevmode\nobreak\ \times\left(2+\frac{\rho\pi}{2\log m_{1}}+\frac{\rho\pi}{2\log m_{2}}\right)\Omega\log b_{0}\end{split}$
to obtain (32).
Finally, we must settle the case (C). We see that
(70) $(R_{i+1,1}+1)(R_{i+2,1}+1)<(1+\epsilon_{i1})(\chi_{1}C_{1})^{4/3}a_{i}^{2}a_{i+1}a_{i+2}\log^{2}b_{0}$
for $i=1,2,3$, where $\epsilon_{i1}=1/R_{i+1,1}+1/R_{i+2,1}+1/(R_{i+1,1}R_{i+2,1})$. Since
(71) $M\geq\chi V>\chi^{2}K>\chi_{1}^{2}\left(1-\frac{1}{K}\right)(C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0})$
and
(72) $\max\{R_{i+1,1},R_{i+2,1}\}\leq(\chi_{1}C_{1})^{2/3}a_{i}\max\{a_{i+1},a_{i+2}\}\log b_{0},$
we have
(73) $M-\max\{R_{i+1,1},R_{i+2,1}\}>\chi_{1}^{2}(1-\epsilon_{i2})C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0},$
where
(74) $\epsilon_{i2}=\frac{\chi^{2/3}}{C_{1}^{1/3}\min\{a_{i+1},a_{i+2}\}\log B_{0}}.$
Hence, we obtain
(75) $\left\lvert u_{i}\right\rvert\leq\frac{(R_{i+1,1}+1)(R_{i+2,1}+1)}{M-\max\{R_{i+1,1},R_{i+2,1}\}}<(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}a_{i}$
for each $i=1,2,3$, where $\epsilon=\max_{1\leq i\leq 3}(1+\epsilon_{i1})/(1-\epsilon_{i2}-1/K)-1$.
We begin with the case $u_{1}\neq 0$. We use
(76) $\begin{split}u_{1}\Lambda=&\leavevmode\nobreak\ u_{1}(b_{1}\log\alpha_{1}+b_{2}\log\alpha_{2}-b_{3}\log\alpha_{3})\\ =&\leavevmode\nobreak\ b_{2}(u_{1}\log\alpha_{2}-u_{2}\log\alpha_{1})-b_{3}(u_{3}\log\alpha_{1}+u_{1}\log\alpha_{3})\\ =&\leavevmode\nobreak\ b_{2}\log\alpha_{1}^{\prime}+b_{3}\log\alpha_{2}^{\prime},\end{split}$
where we put
(77) $\log\alpha_{1}^{\prime}=u_{1}\log\alpha_{2}-u_{2}\log\alpha_{1},\log\alpha_{2}^{\prime}=u_{3}\log\alpha_{1}+u_{1}\log\alpha_{3}.$
We recall that $\alpha_{1}={\sqrt{-1}}$, $\alpha_{2}=\eta_{1}$, $\alpha_{3}=\eta_{2}$ to obtain
(78) $\begin{split}&\rho_{1}\left\lvert\log\alpha_{i}^{\prime}\right\rvert+2\operatorname*{h}(\alpha_{i}^{\prime})<u_{1}\left(\log m_{i}+\frac{\rho_{1}\pi}{2}\right)+\frac{u_{i+1}\rho_{1}\pi}{2}\\ &<(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}\left(\left(\log m_{i}+\frac{\rho_{1}\pi}{2}\right)\rho+\left(\log m_{i}+\frac{\rho\pi}{2}\right)\frac{\rho_{1}\pi}{2}\right)\\ &=(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}((\rho+\rho_{1})\log m_{i}+\rho\rho_{1}\pi)\pi/2.\end{split}$
So that we can take
(79) $a_{i}^{\prime}=(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}((\rho+\rho_{1})\log m_{i}+\rho\rho_{1}\pi)\pi/2$
for $i=1,2$ and therefore
(80) ${b^{\prime\prime}}=\frac{b_{2}}{a_{2}^{\prime}}+\frac{b_{3}}{a_{1}^{\prime}}<\frac{2\chi_{1}^{2/3}}{\pi(\rho+\rho_{1})C_{1}^{1/3}}\left(\frac{e_{1}}{\log m_{2}}+\frac{e_{2}}{\log m_{1}}\right).$
Now we apply Lemma 3.2 with $\rho_{1}$ taken from Table 2 and $\mu=0.61$ to obain
(81) $-\log\left\lvert\Lambda\right\rvert<A_{1}\Omega\log^{2}{b^{\prime\prime}}<A_{1}\Omega\log^{2}(A_{2}^{{\prime\prime}}E),$
where $A_{2}^{\prime\prime}=2\chi_{1}^{2/3}/(\pi(\rho+\rho_{1})C_{1}^{1/3})$. Now we easily obtain (31) observing that $A_{2}^{\prime\prime}<A_{2}<A_{2}\sqrt{\log m_{1}}$.
Finally, if $u_{1}=0$, then $u_{2}b_{2}+u_{3}b_{3}=0$ and therefore we can take $u_{2}=b_{3}^{\prime\prime}=e_{2}/d_{2}$ and $u_{3}=-b_{2}^{\prime}=-e_{1}/d_{2}$. Hence,
(82) $\left\lvert b_{2}^{\prime}\right\rvert\leq\frac{(R_{1,1}+1)(R_{2,1}+1)}{M-\max\{R_{1,1},R_{2,1}\}}<(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}a_{3}$
and
(83) $\left\lvert b_{3}^{\prime\prime}\right\rvert\leq\frac{(R_{1,1}+1)(R_{3,1}+1)}{M-\max\{R_{1,1},R_{3,1}\}}<(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}a_{2}.$
Lemma 2.2 immediately gives that $d_{2}=\gcd(e_{1},e_{2})\leq 4$ and
(84) $e_{1}<4(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}\left(\frac{\rho\pi}{2}+\log m_{2}\right)$
and
(85) $e_{2}<4(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}\left(\frac{\rho\pi}{2}+\log m_{1}\right).$
Hence, we obtain
(86) $E<4(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}\left(2+\frac{\rho\pi}{2\log m_{1}}+\frac{\rho\pi}{2\log m_{2}}\right)<A_{0}.$
This completes the proof of the theorem.
For example, in the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$, we have $K\geq 636133$, $L\geq 42$, and $R_{i}\geq 33242$ for $i=1,2,3$. Hence, we obtain $R_{1}R_{2}R_{3}<18.58295K^{2}L$, $g<0.2455157$, $C_{3}<31.908855$, and therefore the right of (60) is at most $46.93798$. On the other hand, we have
(87) $C_{0}C_{1}\left(1-\frac{1}{K}\right)\left(1-\frac{1}{L}-\frac{1}{3L^{2}}\right)>46.93857$
and we can confirm (60).
Now we can apply Lemma 3.1. In the case (A), proceeding as above, we are led to
(88) $-\log\left\lvert\Lambda\right\rvert<5381.263\Omega\log^{2}(1.985977E\sqrt{\log m_{1}}).$
In the case (B), we have from (69) that
(89) $e_{1}\log m_{1}+e_{2}\log m_{2}<109.7095\Omega\log b_{0}<219.419\Omega\log(1.985977E\sqrt{\log m_{1}}).$
In the case (C), applying Lemma 3.2 with $\rho_{1}=6.2$ and $\mu=0.61$ and observing that
(90) ${b^{\prime\prime}}<\frac{2\chi_{1}^{2/3}}{C_{1}^{1/3}(\rho+\rho_{1})\pi}\left(\frac{e_{1}}{\log m_{1}}+\frac{e_{2}}{\log m_{2}}\right)<0.014249E,$
we obtain
(91) $-\log\left\lvert\Lambda\right\rvert<5381.263\Omega\log^{2}{b^{\prime\prime}}<5381.263\Omega\log^{2}(0.014249E)$
or $E<A_{0}$. These estimates give the theorem in the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$. ∎
Table 2. Other constants used in the proof of Theorem 3.3
$m_{1}$ $\tilde{m}_{2}$ $\chi_{1}$ $C_{0}$ $C_{1}$ $\rho_{1}$ $K_{0}$ $L_{0}$ $R_{0}$
$5$ $2000000005$ $0.46$ $3.2$ $15.499$ $6.0$ $549359$ $42$ $26154$
$13$ $769230773$ $0.46$ $3.3$ $14.320$ $6.1$ $569527$ $43$ $28789$
$17$ $588235297$ $0.50$ $3.3$ $11.653$ $6.8$ $583575$ $40$ $30786$
$29$ $344827589$ $0.48$ $3.3$ $12.650$ $6.5$ $596786$ $41$ $31642$
$37$ $270270277$ $0.53$ $3.4$ $9.571$ $7.3$ $602446$ $39$ $33973$
$41$ $243902441$ $0.47$ $3.3$ $13.519$ $6.2$ $608540$ $42$ $32359$
$5$ $10^{16}$ $0.48$ $3.3$ $13.487$ $6.3$ $822747$ $42$ $26520$
$13$ $10^{16}$ $0.49$ $3.3$ $12.744$ $6.4$ $907924$ $41$ $29588$
$17$ $10^{16}$ $0.54$ $3.3$ $10.207$ $7.1$ $943815$ $38$ $32235$
$29$ $10^{16}$ $0.52$ $3.3$ $11.000$ $6.9$ $979765$ $39$ $32936$
$37$ $10^{16}$ $0.55$ $3.4$ $9.013$ $7.7$ $984944$ $38$ $34781$
$41$ $10^{16}$ $0.51$ $3.3$ $11.673$ $6.6$ $1008482$ $40$ $33538$
$5$ $10^{24}$ $0.48$ $3.3$ $13.435$ $6.4$ $1163620$ $42$ $26532$
$13$ $10^{24}$ $0.49$ $3.3$ $12.708$ $6.5$ $1282306$ $41$ $29594$
$17$ $10^{24}$ $0.54$ $3.4$ $9.648$ $7.3$ $1290303$ $39$ $32099$
$29$ $10^{24}$ $0.52$ $3.4$ $10.388$ $7.1$ $1340720$ $40$ $32759$
$37$ $10^{24}$ $0.55$ $3.5$ $8.533$ $7.9$ $1345753$ $39$ $34693$
$41$ $10^{24}$ $0.51$ $3.3$ $11.658$ $6.7$ $1421332$ $40$ $33543$
$5$ $e^{40000}$ $0.51$ $3.3$ $11.847$ $6.8$ $739807372$ $40$ $27513$
$13$ $e^{40000}$ $0.51$ $3.4$ $11.030$ $7.0$ $786646356$ $41$ $30127$
$17$ $e^{40000}$ $0.59$ $3.4$ $8.245$ $7.9$ $817780793$ $37$ $34383$
$29$ $e^{40000}$ $0.57$ $3.4$ $8.813$ $7.6$ $846771304$ $38$ $34859$
$37$ $e^{40000}$ $0.64$ $3.4$ $7.019$ $8.7$ $888581986$ $35$ $39397$
$41$ $e^{40000}$ $0.57$ $3.3$ $9.513$ $7.3$ $898313919$ $37$ $35840$
$\tilde{m}_{1}$ $\tilde{m}_{2}$ $\chi_{1}$ $C_{0}$ $C_{1}$ $\rho_{1}$ $K_{0}$ $L_{0}$ $R_{0}$
$53$addstackgap[.5]$m_{1}^{*}$ $0.46$ $3.2$ $15.029$ $6.2$ $636133$ $42$ $33242$
$53$ $10^{24}$ $0.49$ $3.3$ $12.550$ $6.6$ $1452862$ $41$ $33647$
$53$ $10^{32}$ $0.51$ $3.3$ $11.688$ $6.7$ $1875132$ $40$ $34263$
$53$ $10^{34}$ $0.51$ $3.3$ $11.683$ $6.8$ $1980530$ $40$ $34264$
$53$ $e^{2000}$ $0.53$ $3.4$ $10.254$ $7.1$ $44504307$ $40$ $34762$
$53$ $e^{40000}$ $0.53$ $3.3$ $10.813$ $7.1$ $913052819$ $39$ $34978$
$10^{24}$ $10^{24}$ $0.64$ $3.5$ $6.286$ $9.1$ $6409655$ $35$ $176815$
$5000$ $e^{2000}$ $0.59$ $3.4$ $8.211$ $7.9$ $59497348$ $37$ $49865$
$5000$ $e^{40000}$ $0.59$ $3.4$ $8.203$ $7.9$ $1184137409$ $37$ $49870$
$10^{7}$ $e^{40000}$ $0.63$ $3.5$ $6.784$ $8.7$ $1604506209$ $36$ $72224$
$10^{14}$ $e^{40000}$ $0.69$ $3.5$ $5.745$ $9.7$ $2520730598$ $34$ $120077$
Table 3. Other constants used in the proof of Theorem 3.3
$m_{1}$ $\tilde{m}_{2}$ $g_{0}$ $C_{3,0}$ $X_{0}$ $Y_{0}$ $A_{2}^{\prime}$
$5$ $2000000005$ $0.2455091$ $32.907152$ $48.406238$ $48.406463$ $4.593026$
$13$ $769230773$ $0.2454815$ $31.827636$ $46.147713$ $46.148423$ $4.481626$
$17$ $588235297$ $0.2457133$ $25.832053$ $37.485114$ $37.485451$ $4.173153$
$29$ $344827589$ $0.2456108$ $28.067845$ $40.717910$ $40.718483$ $4.291573$
$37$ $270270277$ $0.2458429$ $22.128118$ $31.699167$ $31.699820$ $3.911263$
$41$ $243902441$ $0.2455453$ $30.022687$ $43.541755$ $43.541991$ $4.392677$
$5$ $10^{16}$ $0.2455813$ $29.951582$ $43.438633$ $43.438944$ $4.393314$
$13$ $10^{16}$ $0.2456439$ $28.276664$ $41.020708$ $41.021078$ $4.307492$
$17$ $10^{16}$ $0.2459050$ $22.581641$ $32.788675$ $32.788892$ $3.991684$
$29$ $10^{16}$ $0.2458083$ $24.360773$ $35.360809$ $35.361239$ $4.094194$
$37$ $10^{16}$ $0.2459353$ $20.816751$ $29.829780$ $29.830669$ $3.833814$
$41$ $10^{16}$ $0.2457464$ $25.876539$ $37.549576$ $37.549815$ $4.179917$
$5$ $10^{24}$ $0.2455822$ $29.835631$ $43.270668$ $43.271477$ $4.387128$
$13$ $10^{24}$ $0.2456445$ $28.197090$ $40.905122$ $40.905213$ $4.303060$
$17$ $10^{24}$ $0.2458731$ $22.305982$ $31.954006$ $31.954878$ $3.926510$
$29$ $10^{24}$ $0.2457767$ $24.040295$ $34.428320$ $34.428836$ $4.026041$
$37$ $10^{24}$ $0.2459042$ $20.559644$ $29.092665$ $29.093151$ $3.772791$
$41$ $10^{24}$ $0.2457467$ $25.843209$ $37.501236$ $37.501573$ $4.177959$
$5$ $e^{40000}$ $0.2457432$ $26.261907$ $38.108908$ $38.109577$ $4.202487$
$13$ $e^{40000}$ $0.2457150$ $25.549364$ $36.579365$ $36.579880$ $4.111220$
$17$ $e^{40000}$ $0.2460916$ $19.023386$ $27.268387$ $27.268525$ $3.725578$
$29$ $e^{40000}$ $0.2460016$ $20.355247$ $29.168248$ $29.168751$ $3.810001$
$37$ $e^{40000}$ $0.2463000$ $16.156984$ $23.175985$ $23.176260$ $3.529766$
$41$ $e^{40000}$ $0.2460292$ $21.023538$ $30.536539$ $30.536799$ $3.901703$
$\tilde{m}_{1}$ $\tilde{m}_{2}$ $g_{0}$ $C_{3,0}$ $X_{0}$ $Y_{0}$ $A_{2}^{\prime}$
$53$addstackgap[.5]$m_{1}^{*}$ $0.2455157$ $31.908855$ $46.937927$ $46.938571$ $4.541758$
$53$ $10^{24}$ $0.2456472$ $27.846484$ $40.396512$ $40.396637$ $4.283496$
$53$ $10^{32}$ $0.2457461$ $25.909877$ $37.597898$ $37.598084$ $4.181873$
$53$ $10^{34}$ $0.2457462$ $25.898758$ $37.581778$ $37.582001$ $4.181221$
$53$ $e^{2000}$ $0.2458120$ $23.730227$ $33.984228$ $33.984745$ $4.010986$
$53$ $e^{40000}$ $0.2458444$ $23.947111$ $34.760112$ $34.760133$ $4.072742$
$10^{24}$ $10^{24}$ $0.2462967$ $15.080266$ $21.366270$ $21.366409$ $3.403729$
$5000$ $e^{2000}$ $0.2460924$ $18.944866$ $27.155867$ $27.156077$ $3.719988$
$5000$ $e^{40000}$ $0.2460926$ $18.926542$ $27.129543$ $27.129619$ $3.718673$
$10^{7}$ $e^{40000}$ $0.2462366$ $16.294300$ $23.078301$ $23.078337$ $3.496385$
$10^{14}$ $e^{40000}$ $0.2464633$ $13.764710$ $19.509711$ $19.510304$ $3.309117$
Furthermore, we need a lower bound for linear forms in one logarithm and $\pi{\sqrt{-1}}/2$.
Lemma 3.4.
Let
(92) $\Lambda_{1}=b_{2}\log\alpha-\frac{b_{1}\pi{\sqrt{-1}}}{2},$
where $b_{1},b_{2}$ are positive integers, $\alpha$ is an complex algebraic number of absolute value one but not a root of unity and $\log\alpha$ takes its value such that $\left\lvert\operatorname*{Im}\log\alpha\right\rvert<\pi/2$. Put
(93) $\begin{split}{b^{\prime}}=&\frac{b_{1}}{9.05\pi+2D\operatorname*{h}(\alpha)}+\frac{b_{2}}{9.05\pi},\\ D=&[{\mathbb{Q}}(\alpha):{\mathbb{Q}}]/2,\\ a=&9.05\pi+2D\operatorname*{h}(\alpha),\\ H=&\max\{1000,D,D(\log{b^{\prime}}+2.96)+0.01\}.\end{split}$
Then,
(94) $\log\left\lvert\Lambda_{1}\right\rvert>-1.8793aH^{2}.$
Proof.
This follows from Theorem 1 of [6] with $D=1$ by setting $\mu=0.59$, $\rho_{1}=18.1$, $\alpha_{1}=i$, $\alpha_{2}=\alpha$, $a_{1}=\rho\pi/2$, $a_{2}=2D\operatorname*{h}(\alpha)+(\rho\pi/2)$, $L=\left\lfloor H+\sqrt{H^{2}+(1/4)}+(1/2)\right\rfloor$, $k=(v_{1}(L)+\sqrt{v_{1}(L)^{2}+4v_{0}(L)v_{2}(L)})^{2}/(2v_{2}(L))^{2}$ to be the positive real number such that $\sqrt{k}$ satisfies the quadratic equation $v_{2}(L)k-v_{1}(L)\sqrt{k}-v_{0}(L)=0$, where
$v_{0}(x)=\frac{1}{4a_{1}}+\frac{4}{3a_{2}}+\frac{x}{12a_{1}},v_{1}(x)=\frac{x}{3},v_{2}(x)=\lambda(x-H),$
and $K=1+\left\lfloor kLa_{1}a_{2}\right\rfloor$, $R_{1}=4$, $S_{1}=\left\lfloor(L+3)/4\right\rfloor$, $R_{2}=1+\left\lfloor\sqrt{(K-1)La_{2}/a_{1}}\right\rfloor$, $S_{2}=1+\left\lfloor\sqrt{(K-1)La_{1}/a_{2}}\right\rfloor$.
Then, we can confirm that $0.2790.2432<\sqrt{k}<$ 0.279, $H>7.5$, $15\leq L<0.92h$, and $K>700$ and then the condition (2) of Theorem 1 of [6]. Moreover, we can also confirm that $R_{2}-1<2a_{2}H^{2}<b_{1}$ and $S_{2}-1<2a_{1}H^{2}<b_{2}$. Hence the condition (1) of Theorem 1 of [6] also holds. Then, confirming that $\sqrt{k}L<0.19665H$, Theorem 1 of [6] yields that $\log\left\lvert\Lambda_{1}\max\{LSe^{LS\left\lvert\Lambda_{1}\right\rvert/(2b_{2})}/(2b_{2}),LRe^{LR\left\lvert\Lambda_{1}\right\rvert/(2b_{1})}/(2b_{1})\}\right\rvert>-1.8792aH^{2}$ and then $\log\left\lvert\Lambda_{1}\right\rvert>-1.8793aH^{2}$. For the detail of the proof, see [21]. ∎
4. Upper bounds for exponents
In this section, we shall prove upper bounds for exponents in (14) or, equivalently, (15).
Theorem 4.1.
Let $\eta_{1}$ and $\eta_{2}$ be Gaussian integers with $-\pi/2<\arg\eta_{i}/\bar{\eta}_{i}<\pi/2$ such that $m_{1}=\eta_{1}\bar{\eta}_{1}$ and $m_{2}=\eta_{2}\bar{\eta}_{2}$ satisfy the condition of Theorem 3.3. If $x,e_{1},e_{2}$ are nonnegative integers such that $x^{2}+1=m_{1}^{e_{1}}m_{2}^{e_{2}}$ or $x^{2}+1=2m_{1}^{e_{1}}m_{2}^{e_{2}}$, then,
(95) $e_{1}\log m_{1}+e_{2}\log m_{2}\leq A_{4}\Omega\log^{2}(A_{5}\log m_{1}),$
where $A_{5}=4A_{2}^{2}(A_{1}+0.6932/\log^{2}(10000A_{2}\sqrt{\log\tilde{m}_{1}}))^{2}$, $A_{4}$ are constants given in Table 4, and $\Omega=(\log m_{1})(\log m_{2})$ as in the previous section. Moreover, if $m_{1}m_{2}<10^{10}$, then we have
(96) $e_{1}\log m_{1}+e_{2}\log m_{2}<1.7628m_{1}m_{2}.$
Table 4. Constants in (95)
$m_{1}$ $\tilde{m}_{2}$ $A_{4}$
$5$ $2000000005$ $7959.719$
$13$ $769230773$ $6815.554$
$17$ $588235297$ $3981.632$
$29$ $344827589$ $4533.458$
$37$ $270270277$ $3127.262$
$41$ $243902441$ $5922.421$
$5$ $10^{16}$ $7192.899$
$13$ $10^{16}$ $6087.270$
$17$ $10^{16}$ $3510.530$
$29$ $10^{16}$ $3979.106$
$37$ $10^{16}$ $2698.486$
$41$ $10^{16}$ $5206.488$
$5$ $10^{24}$ $6811.925$
$13$ $10^{24}$ $5757.606$
$17$ $10^{24}$ $3305.365$
$29$ $10^{24}$ $3754.144$
$37$ $10^{24}$ $2533.584$
$41$ $10^{24}$ $4914.868$
$5$ $e^{40000}$ $6033.728$
$13$ $e^{40000}$ $5083.203$
$17$ $e^{40000}$ $2875.896$
$29$ $e^{40000}$ $3281.774$
$37$ $e^{40000}$ $2176.067$
$41$ $e^{40000}$ $4350.349$
$\tilde{m}_{1}$ $\tilde{m}_{2}$ $A_{4}$
$53$addstackgap[.5]$m_{1}^{*}$ $6240.476$
$53$ $10^{24}$ $5167.476$
$53$ $10^{32}$ $5024.642$
$53$ $10^{34}$ $4997.394$
$53$ $e^{2000}$ $4579.094$
$53$ $e^{40000}$ $4555.827$
$10^{24}$ $10^{24}$ $1752.029$
$5000$ $e^{2000}$ $2889.083$
$5000$ $e^{40000}$ $2876.189$
$10^{7}$ $e^{40000}$ $2131.592$
$10^{14}$ $e^{40000}$ $1675.249$
Proof.
Assume that $m_{1}$ and $m_{2}$ satisfy the condition of Theorem 3.3 and put $\tilde{m}_{1}$, $\tilde{m}_{2}$, $\Omega$, and $E$ as in Theorem 3.3. We may assume that $e_{1}e_{2}\neq 0$ since $e_{i}=0$ implies that $e_{3-i}=1,2$ or $4$ by Lemma 2.2, so that we can put $\alpha=(e_{2}\log m_{2})/(e_{1}\log m_{1})$. Moreover, we may assume that $E\geq 10000$ since otherwise we have $e_{1}\log m_{1}+e_{2}\log m_{2}<10000\Omega<A_{4}\Omega\log^{2}A_{5}$.
We can decompose $m_{i}=\eta_{i}\bar{\eta}_{i}$ in a way such that $-\pi/4<\arg\eta^{\prime}<\pi/4$. We put $\xi_{i}=\eta_{i}/\bar{\eta}_{i}$ and write $\theta_{i}=\left\lvert\arg\xi_{i}\right\rvert=\left\lvert\log\xi_{i}\right\rvert$, so that $\theta_{i}<\pi/2$.
Now $\Lambda=\log[(x+{\sqrt{-1}})/(x-{\sqrt{-1}})]$ can be represented as a linear form of three logarithms
(97) $\Lambda=\pm e_{1}\log\xi_{1}\pm e_{2}\log\xi_{2}\pm\frac{e_{3}\pi{\sqrt{-1}}}{2}$
for an appropriate integer $e_{3}\geq 0$. Moreover, we can easily see that
(98) $\log\left\lvert\Lambda\right\rvert<-\log(x/2)<-\frac{e_{1}\log m_{1}+e_{2}\log m_{2}}{2}+10^{-9}+\log 2=-\frac{E\Omega}{2}+10^{-9}+\log 2.$
Since $m_{1}$ and $m_{2}$ satisfy the condition of Theorem 3.3, we use this theorem to obtain either (a) (31), (b) (32), or (c) $(A_{2}E)^{2}\log m_{1}\leq B_{0}$.
If (31) holds, then
(99) $\frac{E\Omega}{2}<0.6932+A_{1}\log^{2}(A_{2}E\sqrt{\log m_{1}})\Omega.$
Hence, we have
(100) $A_{2}E\sqrt{\log m_{1}}<2A_{2}\sqrt{\log m_{1}}(0.6932+A_{1}\log^{2}(A_{2}E\sqrt{\log m_{1}}))$
and therefore $A_{2}E\sqrt{\log m_{1}}<\delta\log^{2}(A_{2}E\sqrt{\log m_{1}})$ with $\delta=\sqrt{A_{5}\log m_{1}}$.
Let $c_{5}$ be the unique positive number satisfying
(101) $c_{5}=\left(1+\frac{2\log\log(\sqrt{A_{5}\log\tilde{m}_{1}})+\log c_{5}}{\log(\sqrt{A_{5}\log\tilde{m}_{1}})}\right)^{2}.$
Since $\delta\geq A_{5}\sqrt{\log\tilde{m}_{1}}$, we have
(102) $c_{5}\geq\left(1+\frac{2\log\log\delta+\log c_{5}}{\log\delta}\right)^{2}.$
Now, if $X>c_{5}\delta\log^{2}\delta$, then
(103) $\frac{X}{\log^{2}X}>\frac{c_{5}\delta\log^{2}\delta}{(\log\delta+2\log\log\delta+\log c_{5})^{2}}=\frac{c_{5}\delta}{\left(1+\frac{2\log\log\delta+\log c_{5}}{\log\delta}\right)^{2}}\geq\delta.$
Hence, we must have $A_{2}E\sqrt{\log m_{1}}<c_{5}\delta\log^{2}\delta$ and
(104) $E<\frac{c_{5}}{2}(A_{1}+0.6932/\log^{2}(10000A_{2}\sqrt{\log\tilde{m}_{1}}))\log^{2}(A_{5}\log m_{1}).$
Since $A_{4}\geq c_{5}(A_{1}+0.6932/\log^{2}(10000A_{2}\sqrt{\log\tilde{m}_{1}}))/2$, (95) holds.
If (32) holds, then we have
(105) $A_{2}E\sqrt{\log m_{1}}<A_{2}A_{3}\log(A_{2}E\sqrt{\log m_{1}})\sqrt{\log m_{1}}$
and, proceeding like above, we have $E<A_{4}\log^{2}(A_{5}\log m_{1})$, which immediately yields (95).
If $(A_{2}E)^{2}\log m_{1}\leq B_{0}$, then we immediately have $A_{2}E<A_{2}E\sqrt{\log m_{1}}<\sqrt{B_{0}}$ and $E<\sqrt{B_{0}}/A_{2}$. But we can confirm that $\sqrt{B_{0}}/A_{2}<A_{4}\log^{2}A_{5}$ in any case. Hence, we obtain (95).
Finally, we assume that $m_{1}m_{2}<10^{10}$. We observe that there exist two squarefree integers $D_{1}$ and $D_{2}$ such that $D_{1}D_{2}$ divides $2m_{1}m_{2}$ and $x^{2}-D_{1}y^{2}=-1$ for some integer $y$ all of whose prime factors divide $D_{2}$. We have $x+y\sqrt{D_{1}}=\epsilon^{n}$ for some integer $n$, where $\epsilon$ is the fundamental unit in ${\mathbb{Q}}(D_{1})$ Now it follows from Proposition 3.4.5 of [4, p. 138] and Proposition 10.3.16 of [5, p. 200] with $D_{1}$ in place of $m$ and $f$ in the latter proposition that
(106) $\log\epsilon<D_{1}^{1/2}((\log D_{1})/2+\log\log D_{1}+2.8)<0.4407D_{1}.$
By Carmichael’s theorem [2] (for a simple proof, see [19]), we see that $n$ divides $p-(D_{1}/p)$ for some prime factor $p$ of $y$ (and therefore $D_{2}$) whenever $n>12$. Hence, we have $n\leq D_{2}+1$ to obtain
(107) $\log(x+y\sqrt{D_{1}})=n\log\epsilon<0.4407D_{1}(D_{2}+1)<0.8814m_{1}m_{2},$
which yields that
(108) $e_{1}\log m_{1}+e_{2}\log m_{2}\leq\log(x^{2}+1)=\log(D_{1}y^{2})<1.7628m_{1}m_{2},$
which completes the proof of the lemma. ∎
5. Proof of the Theorem
Let $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}$ and $r$ be integers with $x_{1},x_{2},x_{3}>1,r\neq 0$ and $\gcd(y_{1},y_{2},y_{3})=1$ satisfying (13) and $m_{1},m_{2},s_{i},k_{i},\ell_{i}(i=1,2,3)$ be corresponding integers. Furthermore, let $\eta_{1},\eta_{2}$ be Gaussian integers satisfying (14) and (15). We write $F=\max k_{i}$, $G=\max\ell_{i}$, and, as in preceding sections, $\Omega=(\log m_{1})(\log m_{2})$. Moreover, we write the quantity $\max\{A,B\}$ as $A\lor B$ for brevity.
We note that, since we have assumed that $\{x_{1},x_{2},x_{3}\}\neq\{2,3,7\}$, (13) is nondegenerate by Lemma 2.1 and therefore we have $y_{1}\mid(\pm k_{2}\ell_{3}\pm k_{3}\ell_{2}),y_{2}\mid(\pm k_{3}\ell_{1}\pm k_{1}\ell_{3})$ and $y_{3}\mid(\pm k_{1}\ell_{2}\pm k_{2}\ell_{1})$ with appropriate choices of signs. Hence, $\left\lvert y_{i}\right\rvert\leq 2FG$ for $i=1,2,3$.
Now we assume that $m_{1}m_{2}\geq 10^{10}$. Then, the former part of Theorem 4.1 yields that, for each $i$,
(109) $k_{i}\leq A_{4}(\log m_{2})\log^{2}(A_{5}\log m_{1}),$
where we observe that $A_{4}^{2}\log^{4}A_{5}>e^{20}>B_{0}/A_{2}^{2}$ in any case, and, similarly,
(110) $\ell_{i}\leq A_{4}(\log m_{1})\log^{2}(A_{5}\log m_{1}).$
Hence, we have
(111) $FG\leq A_{4}^{2}\Omega\log^{4}(A_{5}\log m_{1}).$
We have two cases: I. $m_{2}$ divides $x_{i}^{2}+1$ for $i=1,2,3$, and II. $x_{i}^{2}+1=m_{1}^{k_{i}}$ or $2m_{1}^{k_{i}}$ for some $i$.
Case I. In this case, we have two possibilities (a) $x_{i}^{2}+1=m_{2}^{\ell_{i}}$ or $2m_{2}^{2\ell_{i}}$ for some $i$ and (b) $m_{1}m_{2}$ divides $x_{i}^{2}+1$ for each $i=1,2,3$.
In the case (a), we arrange indices so that $x_{1}^{2}+1=m_{2}^{\ell_{1}}$ or $2m_{2}^{\ell_{1}}$. By Lemma 2.2, we have $\ell_{1}=1,2$, or $(x_{1},m_{2},\ell_{1})=(239,13,4)$. In the latter case, we can replace $(x_{1},\ell_{1})$ by $(5,1)$. Hence, we may assume that $\ell_{1}=1$ or $2$.
Since $x_{1}^{2}+1\geq m_{2}$, we have $x_{1}\geq\sqrt{m_{2}-1}$. Since $x_{2}^{2}+1\equiv x_{3}^{2}+1\equiv 0\ \left(\mathrm{mod}\ m_{2}\right)$, Størmer’s criterion implies that $x_{2}>m_{2}/2$ and $x_{3}>m_{2}$.
Since $\left\lvert y_{2}\right\rvert\leq\left\lvert k_{3}\ell_{1}\right\rvert+\left\lvert k_{1}\ell_{3}\right\rvert\leq 2\left\lvert k_{3}\right\rvert$ and $\left\lvert y_{3}\right\rvert\leq\left\lvert k_{2}\ell_{1}\right\rvert+\left\lvert k_{1}\ell_{2}\right\rvert\leq 2\left\lvert k_{2}\right\rvert$, we must have
(112) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{m_{2}-1}}+\frac{2(2\left\lvert k_{2}\right\rvert+\left\lvert k_{3}\right\rvert)}{m_{2}}>\frac{\pi}{4}$
and therefore
(113) $\frac{2FG}{\sqrt{m_{2}-1}}+\frac{6F}{m_{2}}>\frac{\pi}{4}.$
The assumption that $m_{1}m_{2}>10^{10}$ implies that $m_{2}>10^{5}$ and therefore $(2.000011FG+0.018974F)/\sqrt{m_{2}}>\pi/4$. Hence, we have $m_{2}<(2.5465FG+0.02416F)^{2}$.
In the case (b), we can arrange indices so that $x_{3}>x_{2}>x_{1}$. Since $x_{1}^{2}+1\geq m_{1}m_{2}\geq 5m_{2}$, we have $x_{1}\geq\sqrt{5m_{2}-1}$. Since $x_{2}^{2}+1\equiv x_{3}^{2}+1\equiv 0\ \left(\mathrm{mod}\ m_{2}\right)$, Størmer’s criterion implies that $x_{2}>m_{2}/2$ and $x_{3}>m_{2}$.
It immediately follows from (13) with $r\neq 0$ that
(114) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{5m_{2}-1}}+\frac{2\left\lvert y_{2}\right\rvert+\left\lvert y_{3}\right\rvert}{m_{2}}>\frac{\pi}{4}.$
Since $\left\lvert y_{i}\right\rvert\leq 2FG$ for $i=1,2,3$, we have
(115) $\frac{2FG}{(\sqrt{5-10^{-6}})\sqrt{m_{2}}}>\frac{2FG}{\sqrt{5m_{2}-1}}+\frac{6FG}{m_{2}}>\frac{\pi}{4}.$
and therefore $m_{2}<1.306(FG)^{2}$. Hence, in both subcases, we have $m_{2}<(2.5465FG+0.02416F)^{2}$.
On the other hand, we have
(116) $FG<A_{4}^{2}(\log^{2}m_{2})\log^{4}(A_{5}\log m_{2})$
and
(117) $F<A_{4}(\log m_{2})\log^{2}(A_{5}\log m_{2}).$
If $m_{1}\geq 53$ and $m_{2}>1.342\times 10^{34}$, then, we can take $A_{4}=4997.394$ from Table 4. Moreover, we have $A_{5}<168193563$. Hence, we obtain
(118) $m_{2}<4.0444864\times 10^{15}\log^{8}(168193563\log m_{2})\log^{4}m_{2}.$
However, this cannot hold for $m_{2}>1.342\times 10^{34}$. Hence, we must have $m_{1}<m_{2}<1.342\times 10^{34}$ when $m_{1}\geq 53$.
If $m_{2}\geq 10^{34}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{34})$ again, we have $\left\lvert y_{i}\right\rvert<9.098\times 10^{16}$ and $\log x<8380114833$. If $10^{32}\leq m_{2}<10^{34}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{32})$, we have $\left\lvert y_{i}\right\rvert<9.152\times 10^{16}$ and $\log x<8373165769$. If $10^{24}\leq m_{2}<10^{32}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{24})$, we have $\left\lvert y_{i}\right\rvert<8.752\times 10^{16}$ and $\log x<7706422610$. If $m_{2}<10^{24}$, then, applying Theorem 4.1 again with $\tilde{m}_{1}=53$ and $\tilde{m}_{2}=m_{1}^{*}$, we have $\left\lvert y_{i}\right\rvert<7.799\times 10^{16}$ and $\log x<5456058110$. Hence, in any case, we have $y_{i}<9.152\times 10^{16}$ and $x_{i}<\exp(8380114833)$.
Table 5. Upper bounds for $\log x_{i}$ and $y_{i}\leavevmode\nobreak\ (i=1,2,3)$ (Case I)
$m_{1}$ $m_{2}<$ $\log x_{i}<$ $y_{i}<$
$\geq 53$addstackgap[.5]$1.342\times 10^{34}$ $8380114833$ $9.152\times 10^{16}$
$41$ $7.462\times 10^{30}$ $266070791$ $2.146\times 10^{15}$
$37$ $2.157\times 10^{29}$ $108769740$ $3.647\times 10^{14}$
$29$ $1.448\times 10^{30}$ $185150405$ $9.451\times 10^{14}$
$17$ $4.753\times 10^{29}$ $154361476$ $5.415\times 10^{14}$
$13$ $6.816\times 10^{30}$ $324293344$ $2.051\times 10^{15}$
$5$ $5.074\times 10^{30}$ $396883453$ $1.769\times 10^{15}$
Similarly, if $m_{1}=41,37,29,17,13$, and $5$, then $m_{2}<7.461\times 10^{30}$, $2.146\times 10^{29}$, $1.448\times 10^{30}$, $4.753\times 10^{29}$, $6.813\times 10^{30}$, and $5.072\times 10^{30}$, respectively.
Now, applying Theorem 4.1 again with $\tilde{m}_{2}=m_{1}^{*}$, we have $\left\lvert y_{i}\right\rvert\leq 2FG<2.146\times 10^{15}$ and $\log x_{i}<(\log 2+k_{i}\log m_{1}+\ell_{i}\log m_{2})/2<266065055$, that is, $x_{i}<\exp(6263314097)$ for $m_{1}\leq 41$ (see Table 5). This shows the Theorem in Case I.
Case II. We can arrange the indices so as to have $x_{1}^{2}+1=2^{v_{1}}m_{1}^{k_{1}}$ with $v_{1}=0$ or $1$ and $x_{3}>x_{2}$. By Lemma 2.2, we see that $k_{1}=1$, $2$, or $4$ and $k_{1}=4$ implies $x_{1}=239$ and $m_{1}=13$ like above. If $\ell_{i}=0$ for another index $i>1$, then Lemma 2.2 implies that $\{x_{1},x_{2},x_{3}\}=\{2,3,7\}$. Thus, we must have $\ell_{2},\ell_{3}>0$.
We begin by showing that $m_{1}<M_{1}$ with $M_{1}=1.017\times 10^{24}$. Assume that $m_{2}\geq m_{1}\geq M_{1}$. Now, Theorem 4.1 allows us to take $A_{4}=1752.029$, $A_{5}=10545512.8\cdots$, and $\rho=7.292$.
We shall settle the case $k_{1}\leq 2$, so that we leave the case $k_{1}=4$ for later. Hence, we obtain
(119) $\begin{split}FG<A_{4}^{2}\Omega\log^{4}(10545513\log m_{1}).\end{split}$
Similarly, we obtain
(120) $F<A_{4}(\log m_{2})\log^{2}(10545513\log m_{1})$
and
(121) $G<A_{4}(\log m_{1})\log^{2}(10545513\log m_{1}).$
Hence, we obtain
(122) $\left(\frac{G\log m_{2}}{\log m_{1}}\lor F\right)<A_{4}(\log m_{2})\log^{2}(10545513\log m_{1}).$
Now, if $m_{1}\leq 6.5\times 10^{8}G^{2}$, then, (121) gives $m_{1}<6.5\times 10^{8}A_{4}^{2}(\log^{2}m_{1})\log^{4}(10545513\log m_{1})$ and therefore we must have $m_{1}<1.012\times 10^{24}<M_{1}$. Thus, we may assume that $m_{2}>m_{1}>6.5\times 10^{8}G^{2}$.
We clearly have
(123) $y_{1}\arctan\frac{1}{x_{1}}\pm\ell_{3}k_{1}\arctan\frac{1}{x_{2}}\pm\ell_{2}k_{1}\arctan\frac{1}{x_{3}}=\frac{r\pi}{4}$
with $k$ some integer and $y_{1}=\ell_{2}k_{3}\pm\ell_{3}k_{2}$. We can easily see that $x_{2}\geq\sqrt{m_{1}^{k_{2}}m_{2}^{\ell_{2}}-1}\geq\sqrt{m_{2}-1}$. Moreover, from Størmer’s criterion, we observe that $x_{3}>m_{2}/2$. Since we have just assumed that $m_{2}\geq 6.5\times 10^{8}G^{2}$, we obtain
(124) $\left\lvert\frac{r\pi}{4}-y_{1}\arctan\frac{1}{x_{1}}\right\rvert<\frac{k_{1}\ell_{3}}{x_{2}-1}+\frac{k_{1}\ell_{2}}{x_{3}-1}<\frac{(2+10^{-5})G}{\sqrt{m_{2}}}<7.9\times 10^{-5}.$
Like above, we see that $\arctan(1/x_{1})<1/(x_{1}-1)<1/((1-10^{-9})\sqrt{m_{1}})$ and therefore
(125) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{m_{1}}}>(1-10^{-9})\left(\frac{\pi}{4}-7.9\times 10^{-5}\right)>0.78531.$
Since $k_{2}\log m_{1}+\ell_{2}\log m_{2}\leq\log(x_{2}^{2}+1)$, putting $\alpha=\log(x_{2}^{2}+1)/\log m_{2}\geq 1$, we obtain
(126) $\begin{split}\left\lvert y_{1}\right\rvert\leq&\leavevmode\nobreak\ k_{2}G+\ell_{2}F\leq(\log(x_{2}^{2}+1))\left(\frac{G}{\log m_{1}}\lor\frac{F}{\log m_{2}}\right)\\ =&\leavevmode\nobreak\ \alpha\left(\frac{G\log m_{2}}{\log m_{1}}\lor F\right)<A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})\end{split}$
and therefore
(127) $A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})>0.78531\sqrt{m_{1}}.$
Let
(128) $\Lambda_{1}=y_{1}\log\frac{x_{1}+{\sqrt{-1}}}{x_{1}-{\sqrt{-1}}}-\frac{r\pi{\sqrt{-1}}}{2}.$
Then (124) immediately gives that
(129) $\left\lvert\Lambda_{1}\right\rvert<\frac{(4+10^{-9})G}{m_{1}^{k_{2}/2}m_{2}^{\ell_{2}/2}},$
while Lemma 3.4 gives that
(130) $\log\left\lvert\Lambda_{1}\right\rvert>-1.8793aH^{2},$
where
(131) $\begin{split}a=&9.05\pi+\log m_{1}\\ H=&\max\left\{1000,2.97+\log\left(\frac{\left\lvert r\right\rvert}{9.05\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9.05\pi}\right)\right\}.\end{split}$
It immediately follows from (124) and (126) that
(132) $\left\lvert\frac{r\pi}{4}\right\rvert-7.9\times 10^{-5}\leq\frac{\left\lvert y_{1}\right\rvert}{x_{1}-1}<\frac{A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})}{m_{1}^{1/2}}.$
Hence, we obtain
(133) $\frac{\left\lvert r\right\rvert}{9.05\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9.05\pi}<A_{4}\psi_{3}(m_{1})(\alpha\log m_{2})\log^{2}(10545513\log m_{1}),$
where
(134) $\begin{split}\psi_{3}(m_{1})=&\leavevmode\nobreak\ \frac{4}{(9.05\pi+\log m_{1})\pi m_{1}^{1/2}}+\frac{1}{9.05\pi}\\ &\leavevmode\nobreak\ +\frac{3.16\times 10^{-4}}{\pi(9.05\pi+\log m_{1})A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})}.\end{split}$
Since $m_{1}\geq M_{1}$, we have $0.3774942.97+\log\psi_{3}(m_{1})\leq 2.97+\log\psi_{3}(M_{1})<-$ 0.377494. Hence, if $H>1000$, then
[…]
We observe that the right hand side of this inequality is at least $\log(A_{4}\log M_{1})\log^{2}(10545513\log M_{1})>4000$ to obtain
(136) $H<-0.377494+\log(A_{4}\alpha(\log m_{2})\log^{2}(10545513\log m_{1}))$
whether $H>1000$ or not. With the aid of these inequalities, we deduce from (129) and (130) that
[…]
Recalling that $m_{1},m_{2}\geq M_{1}$ and $\alpha\geq 1$, we must have
(138) $\alpha\log m_{2}<103.5805(\log m_{1})\log^{2}(\alpha\log m_{2})$
and therefore
(139) $\alpha\log m_{2}<268.0433(\log m_{1})(\log(103.5805\log m_{1}))^{2}<1246.539(\log m_{1})(\log\log m_{1})^{2}.$
This implies that
(140) $A_{4}(\alpha\log m_{2})\log^{2}(10545513\alpha\log m_{1})<0.78515\sqrt{m_{1}}$
when $m_{1}\geq M_{1}$, which is incompatible with (127). Thus, we are led to the estimate $m_{1}<M_{1}$.
Now we show that $m_{2}<\exp(191323)$. Assume that $m_{2}\geq\exp(191323)$. If $m_{1}\geq M_{1}^{\prime}:=3.537\times 10^{22}$, then, since $m_{2}>e^{40000}$, we iterate the argument that we have just done with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ to obtain (138) and (139) with $28.85883$, $82.15586$, and $281.5761$ in place of $103.5805$, $268.0433$, and $1246.539$. Eventually, we obtain
(141) $A_{4}(\alpha\log m_{2})\log^{2}(6665783\alpha\log m_{1})<0.785395\sqrt{m_{1}}$
with $A_{4}=1675.249$ instead of (140). On the other hand, if $m_{1}\leq 5\times 10^{14}G^{2}$, then, (121) gives $G<A_{4}(\log^{2}(5\times 10^{14}G^{2}))\log^{4}(6665783\log(2.56\times 10^{10}G^{2}))$ and therefore $G<3\times 10^{12}$, which implies that $m_{2}>e^{40000}>5\times 10^{14}G^{2}$. Hence, in any case, we must have $m_{2}>5\times 10^{14}G^{2}$ to obtain (124) and then (127) with $10^{-7}$ in place of $7.9\times 10^{-5}$ and $0.785398$ in place of $0.78531$, contrary to (141). Hene, we must have $m_{1}<M_{1}^{\prime}$ when $m_{2}\geq e^{40000}$.
However, (137) still holds with $2.97+\log\psi_{3}(m_{1})$ in place of $-0.377494$ and $3.16\times 10^{-4}$ in the definition of $\psi_{3}$ replaced by $4\times 10^{-7}$. If $m_{1}\geq 10^{14}$, then we have $2.97+\log\psi_{3}(m_{1})<-0.377493\cdots$ and (137) still holds but this is incompatible with (127) when $m_{2}\geq\exp(191323)$.
If $10^{7}\leq m_{1}<10^{14}$ and $m_{2}\geq e^{40000}$, then we have $2.97+\log\psi_{3}(m_{1})<-0.37723373\cdots$ and (137) still holds with $-0.3772337\cdots$ in place of $-0.377494$. We proceed as just above and obtain $m_{2}<\exp(144270)$. Similarly, if $5000\leq m_{1}<10^{7}$ and $m_{2}\geq e^{40000}$, then we have (137) with $-0.363734\cdots$ in place of $-0.377494$ and $m_{2}<\exp(106130)$. Similarly, if $500053\leq m_{1}<$ 5000 and $m_{2}\geq e^{40000}$, then we have (137) with $-0.234726\cdots$ in place of $-0.377494$ and $m_{2}<\exp(91594)$.
Moreover, if $5\leq m_{1}\leq 41$ and $m_{2}\geq e^{40000}$, then we obtain (137) with $A_{4}$ taken from Table 4 with $\tilde{m}_{2}=e^{40000}$, which is incompatible with (127) when $m_{2}\geq e^{77847}$.
Now we are in turn to settle the remaining case $k_{1}=4$. Then, we must have $x_{1}=239$ and $m_{1}=13$. By Machin’s formula, we have
(142) $4y_{1}\arctan\frac{1}{5}\pm 4\ell_{3}\arctan\frac{1}{x_{2}}\pm 4\ell_{2}\arctan\frac{1}{x_{3}}=\frac{r^{\prime}\pi}{4}$
for some integer $r^{\prime}$, which may be zero. If $r^{\prime}\neq 0$, then we have a reduced case $k_{1}=1$ with $x_{1}=13$ in place of $x_{1}=239$. If $r^{\prime}=0$, then
(143) $y_{1}\arctan\frac{1}{5}=\pm\ell_{3}\arctan\frac{1}{x_{2}}\pm\ell_{2}\arctan\frac{1}{x_{3}}.$
Hence,
(144) $\left\lvert y_{1}\right\rvert\arctan\frac{1}{5}<\frac{\ell_{2}+\ell_{3}}{x_{2}-1}\leq\frac{2G}{x_{2}-1}$
and therefore $m_{2}<((2G/\arctan(1/5))+1)^{2}+1$. If $m_{2}>10^{10}/13$, then
(145) $G\leq\left\lfloor A_{4}(\log 13)\log^{2}(A_{5}\log 13)\right\rfloor=7701562.$
Hence, we have $m_{2}<6.089\times 10^{15}$.
Hence, we must have $m_{1}<1.017\times 10^{24}$ and $m_{2}<\exp(191323)$. Moreover, if $m_{1}\geq 3.537\times 10^{22}$, then $m_{2}<e^{40000}$. If $m_{1}\geq 10^{14}$ and $m_{2}\geq e^{40000}$, then, Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ yields that $\log x\leq(\log 2+k_{i}\log m_{1}+\ell_{i}\log m_{2})/2<3.217\times 10^{12}$, $x_{i}<\exp(3.217\times 10^{12})$, and $\left\lvert y_{i}\right\rvert\leq 2FG<8.334\times 10^{18}$. Similarly, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})$ according to Table 6 in each case, we can confirm that $x_{i}<\exp(3.217\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<8.334\times 10^{18}$ whenever $m_{1}\geq 53$.
Table 6. Upper bounds for $\log x_{i}$ and $y_{i}\leavevmode\nobreak\ (i=1,2,3)$ (Case II)
$m_{1}$ $m_{2}$ $(\tilde{m}_{1},\tilde{m}_{2})$ $\log m_{2}<$ $\log x_{i}<$ $y_{i}<$
[…]
$[10^{7},10^{14}]$ $\geq e^{40000}$ $(10^{7},e^{40000})$ $144270$ $1.963\times 10^{12}$ $6.625\times 10^{18}$
$[5000,10^{7}]$ $\geq e^{40000}$ $(5000,e^{40000})$ $106130$ $9.900\times 10^{11}$ $4.568\times 10^{18}$
$[53,5000]$ $\geq e^{40000}$ $(53,e^{40000})$ $91595$ $7.602\times 10^{11}$ $5.926\times 10^{18}$
$\geq 5000$ $[e^{2000},e^{40000}]$ $(5000,e^{2000})$ $40000$ $1.448\times 10^{12}$ $7.586\times 10^{18}$
$[53,5000]$ $[e^{2000},e^{40000}]$ $(53,e^{2000})$ $40000$ $1.664\times 10^{11}$ $1.300\times 10^{17}$
$\geq 53$ $[m_{1}^{*},e^{2000}]$ $(53,m_{1}^{*})$ $2000$ $1.798\times 10^{12}$ $2.835\times 10^{18}$
$41$ $\geq e^{40000}$ $(41,e^{40000})$ $77848$ $2.411\times 10^{11}$ $1.608\times 10^{18}$
$37$ $\geq e^{40000}$ $(37,e^{40000})$ $71787$ $8.819\times 10^{10}$ $2.401\times 10^{17}$
$29$ $\geq e^{40000}$ $(29,e^{40000})$ $74759$ $1.461\times 10^{11}$ $6.776\times 10^{17}$
$17$ $\geq e^{40000}$ $(17,e^{40000})$ $72529$ $9.837\times 10^{10}$ $3.768\times 10^{17}$
$13$ $\geq e^{40000}$ $(13,e^{40000})$ $76708$ $1.949\times 10^{11}$ $1.544\times 10^{18}$
$5$ $\geq e^{40000}$ $(5,e^{40000})$ $76142$ $1.432\times 10^{10}$ $1.338\times 10^{18}$
$41$ $[10^{24},e^{40000}]$ $(41,10^{24})$ $40000$ $1.498\times 10^{11}$ $1.208\times 10^{18}$
$37$ $[10^{24},e^{40000}]$ $(37,10^{24})$ $40000$ $6.245\times 10^{10}$ $2.160\times 10^{17}$
$29$ $[10^{24},e^{40000}]$ $(29,10^{24})$ $40000$ $9.574\times 10^{10}$ $5.444\times 10^{17}$
$17$ $[10^{24},e^{40000}]$ $(17,10^{24})$ $40000$ $6.701\times 10^{10}$ $3.170\times 10^{17}$
$13$ $[10^{24},e^{40000}]$ $(13,10^{24})$ $40000$ $1.218\times 10^{11}$ $1.156\times 10^{18}$
$5$ $[10^{24},e^{40000}]$ $(5,10^{24})$ $40000$ $8.975\times 10^{10}$ $1.001\times 10^{18}$
$41$ $[10^{16},10^{24}]$ $(41,10^{16})$ $24\log 10$ $223204573$ $1.943\times 10^{15}$
$37$ $[10^{16},10^{24}]$ $(37,10^{16})$ $24\log 10$ $94449003$ $3.577\times 10^{14}$
$29$ $[10^{16},10^{24}]$ $(29,10^{16})$ $24\log 10$ $143666899$ $8.874\times 10^{14}$
$17$ $[10^{16},10^{24}]$ $(17,10^{16})$ $24\log 10$ $100899200$ $5.202\times 10^{14}$
$13$ $[10^{16},10^{24}]$ $(13,10^{16})$ $24\log 10$ $181014051$ $1.850\times 10^{15}$
$5$ $[10^{16},10^{24}]$ $(5,10^{16})$ $24\log 10$ $133264185$ $1.598\times 10^{14}$
$41$ $[243902441,10^{16}]$ $(41,243902441)$ $16\log 10$ $137892418$ $1.112\times 10^{15}$
$37$ $[270270277,10^{16}]$ $(37,270270277)$ $16\log 10$ $57513230$ $1.990\times 10^{14}$
$29$ $[344827589,10^{16}]$ $(29,344827589)$ $16\log 10$ $88172538$ $5.014\times 10^{14}$
$17$ $[588235297,10^{16}]$ $(17,588235297)$ $16\log 10$ $61714770$ $2.920\times 10^{14}$
$13$ $[769230773,10^{16}]$ $(13,769230773)$ $16\log 10$ $112106482$ $1.064\times 10^{15}$
$5$ $[2000000005,10^{16}]$ $(5,2000000005)$ $16\log 10$ $82656260$ $9.218\times 10^{14}$
[…]
Thus we conclude that $x_{i}<\exp(5.306\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<1.472\times 10^{19}$ in Case II.
Finally, we assume that $m_{1}m_{2}<10^{10}$. It immediately follows from the latter part of Theorem 4.1 that
(146) $k_{i}\log m_{1}+\ell_{i}\log m_{2}<1.7628m_{1}m_{2}\leq 1.7628(10^{10}-3),$
[…]
(147) $\left\lvert y_{i}\right\rvert\leq 2FG<\frac{2(0.8814m_{1}m_{2})^{2}}{\Omega}<\frac{2(8.814\times 10^{9})^{2}}{\log 5\log(10^{10}/5)}<4.508\times 10^{18}.$
This completes the proof of the Theorem.
Acknowledgement
This work was supported by the Research Institute for Mathematical Sciences, a joint usage/research center in Kyoto University.